### v1

# Exploring Pessimism and Optimism Dynamics in Deep Reinforcement Learning

Abstract
Off-policy actor-critic algorithms have shown promise in deep reinforcement learning for continuous control tasks. Their success largely stems from leveraging pessimistic state-action value function updates, which effectively address function approximation errors and improve performance. However, such pessimism can lead to under-exploration, constraining the agent’s ability to explore/refine its policies. Conversely, optimism can counteract under-exploration, but it also carries the risk of excessive risk-taking and poor convergence if not properly balanced. Based on these insights, we introduce Utility Soft Actor-Critic (USAC), a novel framework within the actor-critic paradigm that enables independent control over the degree of pessimism/optimism for both the actor and the critic via interpretable parameters. USAC adapts its exploration strategy based on the uncertainty of critics through a utility function that allows us to balance between pessimism and optimism separately. By going beyond binary choices of optimism and pessimism, USAC represents a significant step towards achieving balance within off-policy actor-critic algorithms. Our experiments across various continuous control problems show that the degree of pessimism or optimism depends on the nature of the task. Furthermore, we demonstrate that USAC can outperform state-of-the-art algorithms for appropriately configured pessimism/optimism parameters.
1 Introduction
Deep reinforcement learning (RL) faces significant challenges when navigating the complex landscape of high-dimensional state spaces and non-linear state-action functions. The deadly triad of function approximation, off-policy learning, and bootstrapping compounds these challenges, leading to instability and poor sample efficiency [19, 27, 33]. One particularly critical consequence is the overestimation bias [29], where the estimated state-action values ($Q$-values) exceed their true counterparts [31, 32]. The overestimation bias is typically addressed through a pessimistic approach that use the minimum between two critics’ estimations [10, 12]. While effective in reducing overestimation, this approach can lead to pessimistic under-exploration [7], which limits the agent’s ability to explore and discover new policies. A natural remedy for pessimistic under-exploration could involve integrating optimism into actors’ exploration; however, the balance between pessimism and optimism is not straightforward, as this balance varies across tasks and evolves during the training process [20].
The core focus of this paper is to explore the dynamics between pessimism and optimism in deep RL, especially within actor critics. To address this dynamic, we introduce the novel framework termed Utility Soft Actor-Critic (USAC). Unlike binary approaches, USAC considers pessimism and optimism as points along a spectrum. Our aim is to address the overestimation bias while avoiding pessimistic under-exploration. In the next section, we explore the current literature on this topic and elaborate on our main contributions.
2 Exploring pessimism and optimism in deep reinforcement learning
The primary focus of this work is to understand the interplay between pessimism and optimism and their impact on overestimation bias and pessimistic under-exploration. In this section, we delve into related work addressing these challenges, exploring various strategies to overcome them. Moreover, we conclude this section by outlining our main contributions.
Addressing overestimation bias with pessimism
A pessimistic approach has proven effective in reducing the overestimation bias [^1] The overestimation bias stems from inherent noise and approximation errors introduced by the function approximations and is further amplified by bootstrapping [29, 31, 32]. – as demonstrated in Twin Delayed DDPG [^2] The Deep Deterministic Policy Gradient (DDPG) algorithm was proposed by Lillicrap et al. [17]. (TD3) [10] – it also introduces challenges related to pessimistic under-exploration [7]. Soft Actor-Critic (SAC) [12] addresses this challenge by integrating an entropy term into the actor’s policy decision-making process, thus making the actor stochastic. By incorporating the entropy, SAC encourages exploration and diversifies the actor’s actions, aiming to alleviate the adverse effects of pessimistic under-exploration. However, despite its effectiveness, entropy regularization alone do not resolve the underlying issues.
Exploring optimism to counter pessimistic under-exploration
An intuitive approach to tackle pessimistic under-exploration is based on a general form of the optimism in the face of uncertainty principle. Ciosek et al. [7] introduced the Optimistic Actor-Critic (OAC) algorithm, which adopts an optimistic approximate upper confidence bound on the $Q$-value function to enhance policy exploration. However, it’s crucial to note that aggressive exploration, as encouraged by optimistic algorithms like OAC, can be risky, particularly in environments where the size of overestimation is not well-understood [20].
Figure 1: Impact of pessimism/optimism levels on critic and actor. Top: Final returns. Bottom: Estimation error (i.e., true - estimation); positive values means underestimation, whereas negative values means overestimation.
Balancing between pessimism and optimism
The balance of pessimism and optimism is nuanced and it varies across different tasks and develops throughout the training process [20]. The Tactical Optimism and Pessimism (TOP) algorithm, introduced by Moskovitz et al. [20], addresses this variability by dynamically balancing pessimism and optimism using multi-armed bandits [6, 16]. However, unlike algorithms such as OAC, which lean towards optimism, TOP learns to adjust the degree of pessimism or optimism based on the prevailing uncertainty. However, it uses the same level of pessimism/optimism for both critic and actor training, which limits its versatility.
USAC and our main contributions
Inspired by these insights, we see pessimism and optimism as points along a spectrum rather than binary opposites. While TOP is rooted in the TD3 algorithm, our approach stems from SAC. We introduce a novel framework called USAC, which quantifies the inherent uncertainty in learning through a utility function. Unlike binary approaches, USAC adapts its exploration strategy based on the critics’ distribution through uncertainty estimates of their $Q$-values. Moreover, USAC permits the fine-tuning of pessimism or optimism levels during critic or actor training. Through this methodology, USAC strives for a balanced pessimism-optimism trade-off that overcomes both overestimation bias and pessimistic under-exploration. Figure 1 illustrates the impact of varying levels of pessimism and optimism on both the critic and actor components. This indicates that optimal levels of pessimism and optimism can effectively address both overestimation (or underestimation) and pessimistic under-exploration, thereby enhancing the overall return. Through empirical evaluations and comparisons with state-of-the-art algorithms, we demonstrate the versatility and robustness of USAC across diverse continuous control tasks, paving the way for more adaptive and efficient learning algorithms in the field of deep RL.
3 Preliminary
Throughout this paper, we denote $\mathcal{P}(\Omega)$ as the set of all probability distributions on $\Omega$ and let $\mathcal{B}(\Omega)$ be the set of bounded functions on $\Omega$.
Markov decision processes (MDPs)
[…]
Policies
Let $\Pi=\{\pi:\mathcal{S}\rightarrow\mathcal{P}(\mathcal{A})\}$ be the set of policies. The interaction between the agent and the MDP $\mathcal{M}$ under some policy $\pi\in\Pi$ progresses iteratively: At each time step $t\in\mathbb{N}$, the agent observes state $s_{t}\in\mathcal{S}$, chooses action $a_{t}\in\mathcal{A}$ based on the policy $a_{t}\sim\pi(\cdot|s_{t})$, receives a (bounded) reward $r_{t}:=r(s_{t},a_{t})$, and transitions to the next state $s_{t+1}\sim p(\cdot|s_{t},a_{t})$. Additionally, we define $p^{\pi}(s^{\prime},a^{\prime}|s,a)=p(s^{\prime}|s,a)\pi(a^{\prime}|s^{\prime})$ as the one-step transition probability from $(s,a)$ to $(s^{\prime},a^{\prime})$.
Maximum entropy RL
The standard goal of RL is to find a policy $\pi$ that maximizes the expected sum of discounted rewards $J(\pi)=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$ with initial state $s_{0}\sim p_{0}$ [27, 3, 28]. To improve exploration, we will consider the more general maximum entropy RL framework [36, 11, 12, 25]. In this framework, one aims not only to maximize the expected (discounted) return but also to maximize the (discounted) entropy of the actions suggested by the stochastic policies. Formally, this is achieved by incorporating an entropy term, tempered by a parameter $\alpha>0$, which governs the relative importance of the return compared to the policy entropy. The maximum entropy objective function is defined as $J_{\alpha}(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}(r_{t}+\alpha\mathcal{H}(\pi(\cdot|s_{t})))\right]$, where the policy entropy term $\mathcal{H}(\pi(\cdot|s))=-\mathbb{E}_{a\sim\pi(\cdot|s)}[\log\pi(a|s)]$.
[…]
$T^{\pi}Q(s,a)=r(s,a)+\gamma\mathbb{E}_{(s^{\prime},a^{\prime})\sim p^{\pi}(\cdot|s,a)}[Q(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})]\,,$ (1)
i.e., $T^{\pi}Q^{\pi}(s,a)=Q^{\pi}(s,a)$ for any $(s,a)\in\mathcal{S}\times\mathcal{A}$.
Actor-critic algorithms
One approach to learning the optimal policy is through the framework of actor-critic algorithms. In these algorithms, an agent simultaneously learns a policy (the actor), responsible for selecting actions that maximize expected return, and an estimation of the state-action value function (the critic), responsible for evaluating the policy’s quality through iterative updates. Formally, the critic $Q$ estimates the state-action value function $Q^{\pi}$ (derived by following policy $\pi$) by
$\operatorname*{arg\,min}_{Q}\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}[(Q(s,a)-y(s,a,r,s^{\prime}))^{2}],$ (2)
where $(s,a,r,s^{\prime})$ is sampled from the replay buffer $\mathcal{D}$, and the critic’s target value is defined as
$y(s,a,r,s^{\prime})=r+\gamma[\bar{Q}(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})],\quad a^{\prime}\sim\pi(\cdot|s^{\prime}),$ (3)
which is built using a target critic $\bar{Q}$ that may differ from the critic $Q$ [18, 19]. The actor learns its policy by solving the following:
$\operatorname*{arg\,max}_{\pi\in\Pi}\mathbb{E}_{a\sim\pi(\cdot|s)}[\tilde{Q}(s,a)-\alpha\log\pi(a|s)],$ (4)
where $\tilde{Q}$ is (again) an estimate of the state-action value function $Q^{\pi}$, which can differ from both the critic $Q$ and the target critic $\bar{Q}$. In scenarios where the agent do a full Bellman backup without approximation errors, the critic $Q$, the target critic $\bar{Q}$, and the critic used to update the actor $\tilde{Q}$ are all identical to $Q^{\pi}$ [27, 12]. However, variations in $Q$, $\bar{Q}$, and $\tilde{Q}$ are often introduced to address the challenges we outlined in Section 2; we will explore some of these variations in more details below.
Exploring pessimism and optimism in actor-critic algorithms
A common strategy to address overestimation involves estimating the state-action value function using twin critics, along with associated target critics that are delayed versions of these critics [17, 18, 19, 10]. Specifically, let $Q_{1},Q_{2}:\mathcal{S}\times\mathcal{A}\rightarrow{\mathbb{R}}$ be the two critics and $\bar{Q}_{1},\bar{Q}_{2}:\mathcal{S}\times\mathcal{A}\rightarrow{\mathbb{R}}$ the two target critics that are delayed versions of the critics. The pessimistic strategy, employed by TD3 [10], SAC [12] and OAC [7], is to use the minimum of the two target critics $\{\bar{Q}_{k}\}_{k=1,2}$ to compute the target value of the critics in (3):
$\bar{Q}(s,a)=\min_{k=1,2}\bar{Q}_{k}(s,a).$ (5)
This target is used to train both critics $Q_{1}$ and $Q_{2}$. Similarly, the same approach is employed to update the actor (in TD3 and SAC). Here, the critic to update the actor, $\tilde{Q}$ in (4), is determined by the minimum of the two critics $\{Q_{k}\}_{k=1,2}$:
$\tilde{Q}(s,a)=\min_{k=1,2}Q_{k}(s,a).$ (6)
Ciosek et al. [7] noted that the pessimistic approach, described in 5 and 6, can lead to pessimistic under-exploration. To prevent this, they proposed embracing the optimism in the face of uncertainty principle in their OAC algorithm. However, this approach can potentially make exploration too aggressive, thereby negatively impacting overall performance [20]. Instead, the balance between pessimism and optimism should be tailored to the problem’s nature. Moskovitz et al. [20] suggested using the following target critic to achieve this balance:
$\bar{Q}(s,a)=\bar{\mu}(s,a)+\beta\bar{\sigma}(s,a),\quad\beta\in{\mathbb{R}},$ (7)
where the mean $\bar{\mu}$ and (unbiased) standard deviation $\bar{\sigma}$ are constructed using the two target critics $\bar{Q}_{1},\bar{Q}_{2}$; $\bar{\mu}(s,a)=\frac{1}{2}(\bar{Q}_{1}(s,a)+\bar{Q}_{2}(s,a))$ and $\bar{\sigma}(s,a)=\sqrt{\sum_{k=1,2}(\bar{Q}_{k}(s,a)-\bar{\mu}(s,a))^{2}}$. The critic to update the actor, $\tilde{Q}$, can be defined in a similar manner from the two critics $\{Q_{k}\}_{k=1,2}$. Moskovitz et al. [20] classify 7 as optimistic when $\beta\geq 0$ and pessimistic when $\beta<0$. In the special case when $\beta=-1/\sqrt{2}$, 7 simplifies to taking the minimum of the two critics – as in 5 and 6. The TOP algorithm tries to adjust the balance between pessimism and optimism by using a multi-armed bandit algorithm to select $\beta$ from the set $\{-1,0\}$; here $\beta=-1$ corresponds to a pessimistic estimate and $\beta=0$ to an optimistic one. Although $\beta$ can vary over time using a bandit algorithm, TOP uses the same $\beta$ for both $\bar{Q}$ and $\tilde{Q}$.
4 Utility Soft Actor-Critic (USAC): framework and algorithms
Unlike the aforementioned algorithms, our framework decouples the levels of optimism and pessimism for the actor and critic. This novel approach allows us to be pessimistic in learning the critic while being optimistic in actor training. By doing so, we address the overestimation bias observed during critic training and overcome the problem of pessimistic under-exploration faced by the actor. Consequently, our approach promotes both a stable critic and an explorative actor simultaneously.
In this section, we establish the theoretical groundwork for exploring the balance between optimism and pessimism. We introduce the Utility Soft Actor-Critic (USAC) framework, which seamlessly integrates the uncertainty of critics into the actor learning process. Our framework guarantees policy improvement, as detailed in Section 4.1, while offering flexibility in navigating the optimism-pessimism trade-off. Moving on to Sections 4.2 and 4.3, we present the USAC algorithm and demonstrate how specific choices within our framework can emulate the optimism and pessimism strategies of algorithms like TD3 [10], SAC [12], OAC [7], and TOP [20]; in Section 5, we will explore these varying levels of optimism and pessimism across five MuJoCo environments. The pseudo-code of USAC is provided in Algorithm 1.
Before opening the entire discussion, we first introduce how to quantify the uncertainty of the $Q$-functions.
Quantifying the uncertainty of the $Q$-value distribution
Let $\mathcal{Q}\in\mathcal{P}(\mathcal{B}(\mathcal{S}\times\mathcal{A}))$ be a distribution over $Q$-functions, serving as the agent’s estimate of the $Q$-functions. The dispersion of the distribution reflects the uncertainty of the learned $Q$-functions. To characterize such a distribution, we use the utility function of $\mathcal{Q}$, defined by
$U^{\mathcal{Q}}_{\lambda}(s,a)=\frac{1}{\lambda}\log\mathbb{E}_{Q\sim\mathcal{Q}}[\exp(\lambda Q(s,a))],\quad\lambda\in{\mathbb{R}},$ (8)
where $U^{\mathcal{Q}}_{\lambda}(s,a)=\mathbb{E}_{Q\sim\mathcal{Q}}[Q(s,a)]$ when $\lambda=0$. This utility function is commonly used for measuring risk in finance, economics, and decision-making under uncertainty [9, 34]. [^3] The utility function in 8 is also referred to as expected utility, exponential utility, exponential risk measure, generalized mean, or entropic risk measure according to the context [8, 9]. It also shares similarities with the utility functions employed in risk-sensitive RL [14, 26, 22] and distributional RL [24, 1, 2]. It’s important to note that we apply this utility function to the $\mathcal{Q}$ distribution rather than the return distribution, as is typically done in risk-sensitive and distributional RL.
In this work, we use 8 as a measure of pessimism or optimism of the $\mathcal{Q}$ distribution: a positive value of $\lambda$ implies an optimistic view, while a negative value indicates a pessimistic view. This interpretation is clear from the Taylor expansion: $U^{\mathcal{Q}}_{\lambda}=\mathbb{E}_{Q\sim\mathcal{Q}}[Q]+\frac{\lambda}{2}\mathbb{V}_{Q\sim\mathcal{Q}}[Q]+\mathcal{O}(\lambda^{2})$. Consequently, the utility function $U^{\mathcal{Q}}_{\lambda}$ spans the entire spectrum of the $\mathcal{Q}$ distribution by varying $\lambda$. Another important property of the utility is that when $\mathcal{Q}$ is a Dirac delta distribution centered at $Q^{\pi}$, the $Q$-function of a given policy $\pi\in\Pi$, the expected utility function $U^{\mathcal{Q}}_{\lambda}$ is equivalent to $Q^{\pi}$ for any finite $\lambda$.
4.1 USAC framework: policy evaluation, improvement and guarantees
In this section, we will explain the policy evaluation and policy improvement steps, along with the corresponding guarantees for our USAC framework.
[…]
$T^{\pi}U^{\mathcal{Q}}_{\lambda}(s,a)-U^{\mathcal{Q}}_{\lambda}(s,a)=r(s,a)+\gamma\mathbb{E}_{(s^{\prime},a^{\prime})\sim p^{\pi}(\cdot|s,a)}\left[U_{\lambda}^{\mathcal{Q}}(s^{\prime},a^{\prime})-\alpha\log(\pi(a^{\prime}|s^{\prime}))\right]-U_{\lambda}^{\mathcal{Q}}(s,a)$
should be diminished. Nevertheless, it fosters exploration during learning by maintaining a distribution over the $Q$-functions, which can guide exploration.
On the other hand, given a $\mathcal{Q}$ distribution associated with the current policy $\pi$, denoted as $\mathcal{Q}^{\pi}$, we update the policy, similar to 4, through the following maximization problem:
$\operatorname*{arg\,max}_{\pi\in\Pi}\mathbb{E}_{a\sim\pi(\cdot|s)}[U_{\lambda}^{\mathcal{Q}^{\pi}}(s,a)-\alpha\log\pi(a|s)].$
Extending the policy iteration process from SAC to utility functions ensures that the guarantees of policy improvement still hold [12, 11].
4.2 USAC algorithm
The USAC learning process follows the general actor-critic algorithm. In particular, we estimate the desired quantities by samples from a replay buffer $(s,a,r,s^{\prime})\sim\mathcal{D}$. As the operator $T^{\pi}$ should bring the utility $U^{\mathcal{Q}}_{\lambda}$ to its fixed-point $Q^{\pi}$, the USAC framework learns the critic $Q$ (an estimate of $Q^{\pi}$) as in 2 but with the critic’s target value defined by
$y(s,a,r,s^{\prime})=r+\gamma[U_{\lambda_{\text{critic}}}^{\bar{\mathcal{Q}}}(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})],\quad a^{\prime}\sim\pi(\cdot|s^{\prime}),$ (9)
which is built using a target critic distribution $\bar{\mathcal{Q}}$ with some utility parameter $\lambda_{\text{critic}}$ chosen specifically for the critic. Similar to 4, the actor learns its policy by
$\operatorname*{arg\,max}_{\pi\in\Pi}\mathbb{E}_{a\sim\pi(\cdot|s)}[U_{\lambda_{\text{actor}}}^{\tilde{\mathcal{Q}}}(s,a)-\alpha\log\pi(a|s)],$ (10)
using $\tilde{\mathcal{Q}}$ as the estimate for $\mathcal{Q}$ and $\lambda_{\text{actor}}$ as the utility parameter for the actor.
Note that although Moskovitz et al. [20] consider an optimism/pessimism parameter that can switch between two values over time, they use the same parameter for both actor and critic training. In contrast, our work decouples the optimism/pessimism levels of the actor and critic. This novel approach allows us to select a pessimistic parameter $\lambda_{\text{critic}}$ for learning the critic while using an optimistic parameter $\lambda_{\text{actor}}$ for training the actor. This strategy effectively addresses the overestimation bias observed during critic training and mitigates the issue of pessimistic under-exploration encountered by the actor. Consequently, our approach promotes a stable critic and an explorative actor simultaneously.
Later in the experiments presented in Section 5, we will explore various combinations of pessimism and optimism levels. Before proceeding, we will ground the algorithm by considering a specific distribution. It’s worth noting that while we focus on the Laplace distribution in this paper, the framework described in this section is applicable to other distributions as well. The specific choice of Laplace is simply because it makes it more tangible to fine-tune the level of pessimism/optimism in the critic and actor, and to compare with other actor-critic algorithms.
4.3 USAC algorithm: Laplace distribution
The utility formulation in 8 enables us to compute the utility for any distribution with a moment-generating function. Let’s consider a specific case where $\mathcal{Q}$ follows a Laplace distribution. The algorithm overview is provided in Algorithm 1. Here, we denote the utility parameter associated with this distribution as $\kappa$, reflecting its characteristics within the Laplace context.
Proposition 1.
Suppose $\mathcal{Q}$ is a Laplace distribution. Then, for any $(s,a)\in\mathcal{S}\times\mathcal{A}$, the utility function of $\mathcal{Q}$ can be expressed as
$U^{\mathcal{Q}}_{\kappa}(s,a)=\mu_{\mathcal{Q}}(s,a)+g(\kappa)\sigma_{\mathcal{Q}}(s,a)\quad\text{for}\quad\kappa\in(-1,1),$ (11)
where $\mu_{\mathcal{Q}}$ and $\sigma_{\mathcal{Q}}$ represent the mean and standard deviation of $\mathcal{Q}$, respectively, and
$g(\kappa)=\begin{cases}\log(1/(1-\kappa^{2}))/\sqrt{2}\kappa&\>\text{for}\>\kappa\in(-1,1)\backslash\{0\},\\ 0&\>\text{for}\>\kappa=0.\end{cases}$ (12)
The proof of Proposition 1 can be found in Appendix A. Additionally, we present a corresponding proposition for Gaussian-distributed $Q$-values in Appendix B. In Proposition 1, we observe that by varying $\kappa$ in the small interval $(-1,1)$, $U^{\mathcal{Q}}_{\kappa}$ spans the entire spectrum of the $\mathcal{Q}$ distribution. Specifically, as $\lim_{\kappa\rightarrow-1}g(\kappa)\rightarrow-\infty$, $\lim_{\kappa\rightarrow 0}g(\kappa)\rightarrow 0$, and $\lim_{\kappa\rightarrow 1}g(\kappa)\rightarrow\infty$.
Now suppose the uncertainty follows a Laplace distribution. Following the algorithmic steps outlined in Section 4.2, we can estimate the utility of the target critic distribution $\bar{\mathcal{Q}}$ by two target critics $\{\bar{Q}_{k}\}_{k=1,2}$, which are delayed versions of the critics $\{Q_{k}\}_{k=1,2}$. The utility function is given by $U^{\bar{\mathcal{Q}}}_{\kappa_{\text{critic}}}=\mu_{\bar{\mathcal{Q}}}+g(\kappa_{\text{critic}})\sigma_{\bar{\mathcal{Q}}}$ with utility parameter $\kappa_{\text{critic}}\in(-1,1)$, where
$\mu_{\bar{\mathcal{Q}}}(s,a)=\frac{1}{2}(\bar{Q}_{1}(s,a)+\bar{Q}_{2}(s,a))\quad\text{and}\quad\sigma_{\bar{\mathcal{Q}}}(s,a)=\frac{1}{2}\lvert\bar{Q}_{1}(s,a)-\bar{Q}_{2}(s,a)\rvert.$ (13)
Hence, the target value to train the critic becomes
$y(s,a,r,s^{\prime})=r+\gamma[\mu_{\bar{\mathcal{Q}}}(s^{\prime},a^{\prime})+g(\kappa_{\text{critic}})\sigma_{\bar{\mathcal{Q}}}(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})],\quad a^{\prime}\sim\pi(\cdot|s^{\prime}).$ (14)
Similarly, the actor learns its policy by
$\operatorname*{arg\,max}_{\pi\in\Pi}\mathbb{E}_{a\sim\pi(\cdot|s)}[\mu_{\tilde{\mathcal{Q}}}(s,a)+g(\kappa_{\text{actor}})\sigma_{\tilde{\mathcal{Q}}}(s,a)-\alpha\log\pi(a|s)],\quad\kappa_{\text{actor}}\in(-1,1),$ (15)
where $\mu_{\tilde{\mathcal{Q}}}$ and $\sigma_{\tilde{\mathcal{Q}}}$ are estimated in similar manner as 13 on $\tilde{\mathcal{Q}}$ using the critics $\{Q_{k}\}_{k=1,2}$. The pseudo-code of this algorithm can be found in Algorithm 1.
We note that a smaller $\kappa$ value indicates more pessimism, while a larger $\kappa$ value indicates more optimism. By choosing $\kappa$ such that $g(\kappa)=-1$ (i.e., $\kappa\approx-0.83$), the utility in 11 simply reduces to the minimum of the two critics. Thus, choosing $\kappa_{\text{critic}}=\kappa_{\text{actor}}\approx-0.83$, we obtain the pessimism choices of SAC [12] and TD3 [10]. On the other hand, by choosing $\kappa_{\text{critic}}\approx-0.83$ and a positive $\kappa_{\text{actor}}$, we have a pessimism critic and an optimistic actor as explored in OAC [7]. Lastly, by letting a multi-armed bandit algorithm choose $\kappa_{\text{critic}}=\kappa_{\text{actor}}\in\{-0.916563,0\}$, we recover the pessimistic/optimistic design suggested by TOP [20]. [^4] To recover the pessimistic case of the TOP algorithm, we can choose $\kappa=-0.916563$, which corresponds to having the function $g$ in 12 equal to $\sqrt{2}$. We note that by choosing $-0.916563$, TOP considers a more pessimistic critic target than SAC, TD3, and OAC.
In Section 5, we will explore various $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ pairs and discuss whether the previous wisdom about the effectiveness of these strategies holds across different environments.
Algorithm 1 USAC with Laplace distribution (Section 4.3)
1:Input: Averaging parameter $\tau\in(0,1)$, learning rates $\eta_{\theta},\eta_{\phi}>0$, mini-batch size $n\in\mathbb{N}$, pessimism/optimism coefficients $\kappa_{\text{actor}},\kappa_{\text{critic}}\in(-1,1)$
2:Initialize: Replay buffer $\mathcal{D}=\varnothing$, initial state $s_{0}\sim p_{0}$, critic $\{\theta_{k}\}_{k=1,2}$, target critic $\{\bar{\theta}_{k}\}_{k=1,2}$ and actor policy $\phi$
3:for each time step do
4:     for each environment step do
5:         $a_{t}\sim\pi_{\phi}(\cdot|s_{t})$triangleright Sample action from the policy $\phi$
6:         $s_{t+1}\sim p(\cdot|s_{t},a_{t})$triangleright Sample transition from the environment
7:         $\mathcal{D}\leftarrow\mathcal{D}\cup(s_{t},a_{t},r_{t},s_{t+1})$triangleright Store transition in the replay pool
8:     end for
9:     for each training step do
10:         $\{(s_{i},a_{i},r_{i},s_{i}^{\prime},a_{i}^{\prime}):(s_{i},a_{i},r_{i},s_{i}^{\prime})\sim\mathcal{D},a_{i}^{\prime}\sim\pi_{\phi}(\cdot|s_{i}^{\prime})\}_{i=1}^{n}$triangleright Sample mini-batch
11:         for $i\in\{1,\dots,n\}$ do $\triangleright$ Compute target critic distribution $\bar{\mathcal{Q}}$
12:              $\mu_{\bar{\mathcal{Q}}}(s_{i},a_{i})\leftarrow\frac{1}{2}(Q_{\bar{\theta}_{1}}(s_{i},a_{i})+Q_{\bar{\theta}_{2}}(s_{i},a_{i}))$
13:              $\sigma_{\bar{\mathcal{Q}}}(s_{i},a_{i})\leftarrow\frac{1}{2}\lvert Q_{\bar{\theta}_{1}}(s_{i},a_{i})-Q_{\bar{\theta}_{2}}(s_{i},a_{i})\rvert$
14:         end for
15:         for each critic, $k\in\{1,2\}$ do
16:              $\theta_{k}\leftarrow\theta_{k}-\eta_{\theta}\nabla_{\theta_{k}}\left\{\frac{1}{n}\sum_{i=1}^{n}\left(Q_{\theta_{k}}(s_{i},a_{i})-y(s_{i},a_{i},r_{i},s_{i}^{\prime})\right)^{2}\right\}$triangleright Update critic
17:             with $y(s_{i},a_{i},r_{i},s_{i}^{\prime})=r_{i}+\gamma[\mu_{\bar{\mathcal{Q}}}(s_{i}^{\prime},a_{i}^{\prime})+g(\kappa_{\text{critic}})\sigma_{\bar{\mathcal{Q}}}(s_{i}^{\prime},a_{i}^{\prime})-\alpha\log\pi_{\phi}(a_{i}^{\prime}|s_{i}^{\prime})]$
18:         end for
19:         for $i\in\{1,\dots,n\}$ do $\triangleright$ Compute actor distribution $\tilde{\mathcal{Q}}$
20:              $\mu_{\tilde{\mathcal{Q}}}(s_{i},a_{i})\leftarrow\frac{1}{2}(Q_{\theta_{1}}(s_{i},a_{i})+Q_{\theta_{2}}(s_{i},a_{i}))$
21:              $\sigma_{\tilde{\mathcal{Q}}}(s_{i},a_{i})\leftarrow\frac{1}{2}\lvert Q_{\theta_{1}}(s_{i},a_{i})-Q_{\theta_{2}}(s_{i},a_{i})\rvert$
22:         end $\triangleright$ Update actor
23:         $\phi\leftarrow\phi+\eta_{\phi}\nabla_{\phi}\left\{\frac{1}{n}\sum_{i=1}^{n}\left(\mu_{\tilde{\mathcal{Q}}}(s_{i},a_{i})+g(\kappa_{\text{actor}})\sigma_{\tilde{\mathcal{Q}}}(s_{i},a_{i})-\alpha\log\pi_{\phi}(a_{i}|s_{i})\right)\right\}$
24:         for each critic, $k\in\{1,2\}$ do
25:              $\bar{\theta}_{k}\leftarrow\tau\theta_{k}+(1-\tau)\bar{\theta}_{k}$triangleright Update target critic
26:         end for
27:     end for
28:end for
5 Experiments
The goal of our experiments is to explore the interplay between critic and actor pessimism/optimism, estimation error (i.e., overestimation/underestimation), and performance. We compare our USAC algorithm with Laplace distribution (Section 4.3) with prior off-policy actor-critic algorithms. Specifically, we consider SAC [12], TD3 [10], OAC [7] and TOP [20] as baselines, as they represent the most relevant algorithms that incorporate optimism and pessimism. For SAC and USAC, we use auto-tuned entropy temperature $\alpha$ [13], while for OAC and TOP, we adopt the best parameters reported in their respective papers.
We conduct our experiments on 5 continuous control environments in the MuJoCo physics engine [30]. Each experiment is repeated across 5 different seeds, each comprising one million time steps, and evaluated over 10 episodes. The reported results represent the average (along with the standard deviation) across these repetitions. Additional details about architectures, baselines, computational cost, environment properties, grid search, hyper-parameters, learning curves, area under learning curves, estimation error plots and their values can be found in Appendix C.
Insights on USAC performance
In Table 1, we present two sets of results for our USAC algorithm: a default and a best pair of parameters $(\kappa_{\text{critic}},\kappa_{\text{actor}})$; we will discuss how these parameters are selected below. These results demonstrate that our USAC algorithm can outperform (or match) all baseline algorithms across all environments. This success reaffirms our hypothesis, as outlined in Section 4, that the levels of pessimism and optimism should vary between the critic and actor, as well as across environments [20]. In particular, our findings reveals that there exist pairs of utility parameters, $(\kappa_{\text{critic}},\kappa_{\text{actor}})$, capable of overcoming overestimation while facilitating exploration, thus enhancing overall performance. In addition, USAC smaller standard deviation indicates an improved model stability during learning. The corresponding learning curves and estimation error curves for these results are presented in Figures 5 and 6 (Appendix C).For the Humanoid-v4 environment, we found that the default parameter pair outperformed the best pair identified through grid search, e.g., see Figure 2. Therefore, we used the default pair for this environment in our results (Table 1). However, we conjecture that by expanding our (limited) grid search, a better pair may exist.
Table 1: Final return on MuJoCo environments trained with 1M time steps, averaged over 5 seeds. The best algorithms are highlighted in bold. $\pm$ corresponds to the standard deviation across repetitions. The default $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ are both $-0.831559$. The best $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ are listed in Table 2.
Environment USAC (ours) SAC TD3 OAC TOP
Default $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ Best $(\kappa_{\text{critic}},\kappa_{\text{actor}})$
Ant-v4 $\boldsymbol{5139\pm 978}$ $\boldsymbol{5158\pm 1186}$ $4756\pm 1411$ $4091\pm 1303$ $4177\pm 1392$ $4334\pm 1276$
HalfCheetah-v4 $\boldsymbol{11024\pm 849}$ $\boldsymbol{11736\pm 317}$ $10763\pm 895$ $10570\pm 1400$ $8684\pm 1678$ $7311\pm 3074$
Hopper-v4 $3194\pm 810$ $\boldsymbol{3442\pm 247}$ $3185\pm 537$ $1986\pm 1154$ $3293\pm 113$ $\boldsymbol{3367\pm 154}$
Humanoid-v4 $\boldsymbol{5602\pm 505}$ $5602\pm 505^{\star}$ $\boldsymbol{5503\pm 373}$ $5149\pm 737$ $5390\pm 234$ $5332\pm 445$
Walker2d-v4 $\boldsymbol{4525\pm 534}$ $\boldsymbol{4530\pm 767}$ $3757\pm 1282$ $4369\pm 601$ $3467\pm 1200$ $4317\pm 631$
Exploring the dynamics of pessimistic and optimistic with USAC
To explore the dynamics of pessimism and optimism in USAC, we conducted a grid search across the 5 continuous control environments in MuJoCo. More details about this grid search can be found in Appendix C. We examined critics ranging from pessimistic to semi-optimistic and actors from pessimistic to optimistic. Specifically, we used $\kappa_{\text{critic}}\in\{-0.831559,-0.5,-0.33\}$ [^5] Recall that $\kappa=-0.831559$ corresponds to $g(\kappa)=1$, simplifying the utility function in 11 to the minimum of the two critics. Therefore, by setting $\kappa_{\text{critic}}\approx-0.83$, we align with the pessimistic choices of SAC [12], TD3 [10], and OAC [7]., and $\kappa_{\text{actor}}\in\{-0.99,-0.5,0,0.5,0.99\}$. The rationale for choosing more pessimistic (to semi-optimistic) $\kappa_{\text{critic}}$ values is to counteract the overestimation bias. Given that USAC stems from the framework of SAC, SAC was included in our figures for comparison.
Table 2: Best $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$ parameters for USAC.
Environment $\kappa_{\text{critic}}$ $\kappa_{\text{actor}}$
Ant-v4 $-0.831559$ $-0.99$
HalfCheetah-v4 $-0.33$ $-0.50$
Hopper-v4 $-0.831559$ $0.50$
Humanoid-v4 $-0.831559$ $-0.831559^{\star}$
Walker2d-v4 $-0.831559$ $0.0$
By comparing the final returns (Figure 2) with the estimation error results (Figure 3) for the different grid combinations, we reaffirm previous findings [10, 7, 20]: i) overestimation (negative estimation error) should always be avoided as it reduces performance, and ii) underestimation (positive estimation error) does not necessarily harm performance. This indicates that a pessimistic approach should be adopted in critic training, whereas in actor training, either a pessimistic or optimistic approach can be suitable depending on the environment. The best combinations of $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$, with respect to final returns, are summarized in Table 2. Additionally, the area under the learning curve for the grid search parameters is illustrated in Figure 4 (Appendix C).
Figure 2: Grid search results for the final return of USAC algorithm with Laplace distribution, averaged over three seeds, each with three evaluation episodes. The grid search explore various $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$ values, using SAC as the baseline.
Figure 3: Grid search results for the estimation error (true - estimation) of USAC algorithm with Laplace distribution, averaged over three seeds, each with three evaluation episodes. The grid search explore various $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$ values, using SAC as the baseline.
6 Discussion
We introduce Utility Soft Actor-Critic (USAC), a novel off-policy actor-critic algorithm. The key idea is to balance the trade-off between pessimism and optimism through incorporating a utility function that captures the uncertainty in the critic due to limited access to the environment. This framework is flexible and can be adapted to different distributions, with our experiments primarily focusing on the Laplace distribution. Unlike OAC [7] and TOP [20], which also explore optimism/pessimism in actor-critic training, USAC provides a policy improvement guarantee.
USAC allows independent control over the optimism and pessimism levels for the actor and critic. This decoupling enables a stable critic, necessary for accurate value function estimation, and an explorative actor, crucial for effective policy learning. We conducted a grid search experiment to explore these levels in training, revealing that being pessimistic in critic training while being either optimistic or pessimistic in actor training can be beneficial. However, it is also essential to consider the specific environment. Our work paves the way for further exploration in optimizing actor-critic algorithms with varying optimism and pessimism levels.
Limitations and future work
USAC’s primary limitation is its reliance on fixed optimism/pessimism parameters. An exciting direction for future research is to develop adaptive schemes where these parameters adjust dynamically throughout the learning process to better respond to the changing dynamics of the environment. There are several potential approaches to integrate automatic tuning. For example, these include: (i) adapting the automatic tuning scheme used for $\alpha$ in SAC [12, 13], (ii) employing bandit-based selection methods [20], or (iii) exploring continuous learning approaches using gradient descent [4].
Another promising direction for future work is to incorporate higher moments of the critic distributions. By considering not only the mean and variance but also skewness, kurtosis, and heavy-tailed distributions, we can better capture the nuances of uncertainty in the critic’s value estimates. This approach could lead to more robust and flexible policy learning, accommodating a wider variety of distributional characteristics. Incorporating these higher-order moments may provide a richer framework for addressing the biases and variances in the value function estimation.
Disclosure of funding
This work was funded by the Novo Nordisk Foundation (NNF21OC0070621) and the Carlsberg Foundation (CF21-0250).
Appendix A Proofs
Proof of Proposition 1.
Suppose $Q(s,a)\sim\textrm{Laplace}(\mu_{\mathcal{Q}}(s,a),b_{\mathcal{Q}}(s,a))$. Then, for any $(s,a)\in\mathcal{S}\times\mathcal{A}$, the expected utility function of $\mathcal{Q}$ on $(s,a)$ can be expressed as
$U^{\mathcal{Q}}_{\lambda}(s,a)=\mu_{\mathcal{Q}}(s,a)-\lambda^{-1}\log(1-\lambda^{2}b_{\mathcal{Q}}^{2}(s,a)),$
with $|\lambda|<1/b_{\mathcal{Q}}(s,a)$. Next, given that the variance of a Laplace distribution is $2b_{\mathcal{Q}}^{2}$, we can rewrite the expected utility function as
$U^{\mathcal{Q}}_{\lambda}(s,a)=\mu_{\mathcal{Q}}(s,a)-\lambda^{-1}\log(1-\lambda^{2}\sigma_{\mathcal{Q}}^{2}(s,a)/2),$
with $|\lambda|<\sqrt{2}/\sigma_{\mathcal{Q}}(s,a)$. Substituting $\lambda=\sqrt{2}\kappa/\sigma_{\mathcal{Q}}(s,a)$ for some $\kappa\in(-1,1)$, we have
$U^{\mathcal{Q}}_{\kappa}(s,a)=\mu_{\mathcal{Q}}(s,a)-\sigma_{\mathcal{Q}}(s,a)\log(1-\kappa^{2})/\sqrt{2}\kappa.$
At last, this expression can be further simplified using the definition of $g(\kappa)$ in 12.
Appendix B USAC under gaussianity
Proposition 2.
Suppose $\mathcal{Q}$ is Gaussian with mean $\mu_{\mathcal{Q}}$ and variance $\sigma_{\mathcal{Q}}^{2}$, i.e., $Q(s,a)\sim\mathcal{N}(\mu_{\mathcal{Q}}(s,a),\sigma_{\mathcal{Q}}^{2}(s,a))$. Then, for any $(s,a)\in\mathcal{S}\times\mathcal{A}$, the utility function of $\mathcal{Q}$ can be expressed as
$U^{\mathcal{Q}}_{\lambda}(s,a)=\mu_{\mathcal{Q}}(s,a)+\lambda\sigma_{\mathcal{Q}}^{2}(s,a)/2\quad\text{for}\quad\lambda\in(-\infty,\infty).$ (16)
Proof of Proposition 2.
The proof follows by similar steps as in the proof of Proposition 1.
Similarly, we can construct a USAC algorithm under Gaussian uncertainty, analogous to our approach under Laplace uncertainty as discussed in Section 4.3. By using the utility function defined in 16, we can incorporate specific $\lambda_{\text{critic}}$ and $\lambda_{\text{actor}}$ parameters for critic and actor training, respectively. This allows us to adjust the levels of optimism and pessimism tailored to the Gaussian-distributed $\mathcal{Q}$ values, thereby facilitating effective learning and exploration in our framework.
Appendix C Experiments details
Table 3: Properties of MuJoCo environments
Environment Action dimension Observation dimension
Ant-v4 8 27
HalfCheetah-v4 6 17
Hopper-v4 3 11
Humanoid-v4 17 376
Walker2d-v4 6 17
Environments
Our experiment is implemented in PyTorch [21, Version $2.1.0$]. The experiments are conducted on 5 continuous control environments in the MuJoCo physics engine [30, 5]. More detailed information about these environments are provided in Table 3.
Table 4: Shared hyper-parameters
Hyper-parameter Value
Evaluation episodes 10
Evaluation frequency Maximum time steps / 100
Discount factor $(\gamma)$ 0.99
$n$-step returns 1 step
Replay ratio 1
Replay buffer size 1,000,000
Maximum time steps 1,000,000
Mini-batch size $(n)$ 256
Actor/critic optimizer Adam [15]
Optimizer learning rates $(\eta_{\phi},\eta_{\theta},\eta_{\alpha})$ 3e-4
Averaging parameter $(\tau)$ 5e-3
Evaluation and seeds
As mentioned in Table 4, we train our algorithms for one million time steps. Each experiment is repeated across 5 different seeds. We evaluate at every $10,000$ time step for ten evaluation episodes. For training, we use seeds 1 to 5, and for evaluation, we use seeds 101 to 105, which correspond to the training repetition number plus 100.
Hyper-parameters
The hyper-parameters and network configurations used in our experiments can be found in Table 4. For SAC and USAC, we learn the entropy temperature $\alpha$ during as outlined in Haarnoja et al. [13]. For OAC [7] and TOP [20], we adopt the best parameters reported in their respective papers.
Architectures and baselines
The architectural specifications for all algorithms are presented in Table 5. Here, $d_{s}$ and $d_{a}$ represent the dimensions of the state and action spaces, respectively. The SquashedGaussian module features a Gaussian head, which uses the first $d_{a}$ inputs for mean calculation and the subsequent $d_{a}$ inputs for variance. This architecture is the standard one used in original implementations of SAC [12], TD3 [10], and OAC [7]. However, while TOP [20] states that it employs a network architecture with $2$ hidden layers, its original implementation code actually considers $3$ hidden layers. [^6] We use the implementation provided by the authors for TOP, https://github.com/tedmoskovitz/TOP, where the "number of random actions" at the start of training updated to $10000$. This discrepancy prompted us to re-run TOP with a synchronized network such that it aligns with the other algorithms. In addition, the critic network in TOP follows the original implementation with a return shape that equals the number of quantiles (N-Quantiles), which is $50$.
Table 5: Architecture details
Actor network Critic network
USAC
SAC
OAC
TD3
TOP
USAC
SAC
OAC
TD3
TOP
Linear($d_{s}$, 256)
ReLU()
Linear($d_{s}$ + $d_{a}$, 256)
ReLU()
Linear(256, 256)
ReLU()
Linear(256, 256)
ReLU()
Linear(256, 2×$\times d_{a}$, $\theta$)
Linear(256, $d_{a}$) Linear(256, 1) Linear(256, N-Quantiles)
SquashedGaussian(2×$\times d_{a}$, $d_{a}$) Tanh()
Grid search of USAC
In order to explore the dynamics of pessimism and optimism in USAC with Laplace distribution (Section 4.3), we conduct a grid search across five continuous control environments in MuJoCo (see e.g., environments specifications above). These grid search results are averaged over 3 seeds (seed 1 to 3), each with 3 evaluation episodes. This grid search investigated various values of $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$, using SAC as the baseline for comparison.
We examined critics ranging from pessimistic to semi-optimistic and actors from pessimistic to optimistic. Specifically, we used $\kappa_{\text{critic}}\in\{-0.831559,-0.5,-0.33\}$ [^7] A $\kappa=-0.831559$ corresponds to $g(\kappa)=1$, which means that the utility function in 11 simplifies into the minimum of the two critics. Hence, by setting $\kappa_{\text{critic}}\approx-0.83$, we align with the pessimistic choices of SAC [12], TD3 [10], and OAC [7]., and $\kappa_{\text{actor}}\in\{-0.99,-0.5,0,0.5,0.99\}$. The rationale for choosing more pessimistic (to semi-optimistic) $\kappa_{\text{critic}}$ values is to counteract the overestimation bias. Given that USAC stems from the framework of SAC, SAC was included in our figures for comparison.
In Figures 2, 3, and 4, we present the results of our grid search, showcasing the final returns, estimation error, and area under the learning curves for various parameter combinations. The estimation error calculated by subtracting the estimated $Q$-value from the true (discounted) reward. Thus, positive values indicate underestimating, while negative value indicate overestimation. These findings provide insights into the dynamics of pessimism and optimism in our USAC algorithm. As mentioned in Section 5, they reveal that: (i) overestimation (negative estimation error) should always be avoided as it reduces performance, and (ii) underestimation (positive estimation error) does not necessarily harm performance. This suggests that a pessimistic approach should be adopted in critic training, while either a pessimistic or optimistic approach can be suitable for actor training, depending on the environment. The best combinations of $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$ in terms of final returns are summarized in Table 2. These optimal pairs highlight the importance of tailoring the levels of pessimism and optimism to the specific requirements of each environment to enhance overall performance.
Figure 4: Grid search results for the area under the learning curve of USAC algorithm with Laplace distribution, averaged over three seeds, each with three evaluation episodes. The grid search explore various $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$ values, using SAC as the baseline.
Learning curve and estimation error curves
The learning curve and estimation error curves of our USAC algorithm and the baseline algorithms can be seen in Figures 5 and 6, which correspond to the experiments summarized in Table 1. In addition, the estimation error values for these experiments are provided inTable 6. For USAC, the legend corresponds to the default case where $\kappa_{\text{critic}}=\kappa_{\text{actor}}=-0.831559$. The USAC (best) legend uses the optimal $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ pairs (with respect to final return) as listed in Table 2.For the Humanoid-v4 environment, we found that the default parameter pair outperformed the best pair identified through grid search, e.g., see Figure 2. Therefore, we used the default pair for this environment in our results (Table 1). However, we conjecture that by expanding our (limited) grid search, a better pair may exist.
Table 6: Estimation error of USAC and baselines.
Environment USAC (ours) SAC TD3 OAC TOP
Default $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ Best $(\kappa_{\text{critic}},\kappa_{\text{actor}})$
Ant-v4 $80.66$ $91.84$ $75.33$ $196.77$ $205.49$ $98.99$
HalfCheetah-v4 $221.21$ $146.07$ $222.93$ $216.54$ $190.28$ $81.97$
Hopper-v4 $12.46$ $18.51$ $12.73$ $24.82$ $25.05$ $11.73$
Humanoid-v4 $40.24$ $40.24^{\star}$ $42.78$ $31.89$ $52.66$ $24.02$
Walker2d-v4 $42.50$ $44.87$ $36.65$ $50.88$ $52.55$ $34.34$
Figure 5: Learning curves of USAC and baselines. Solid curves depict the average return across evaluation episodes, while the shaded areas represent the standard deviation.
Figure 6: Estimation error curves of USAC and baselines. Solid lines indicate the estimated value function by the critic, while dashed lines show the averaged episodic discounted true return. The shaded areas around the solid and dashed lines represent the standard deviation from the average estimated value and the average true reward, respectively.
Computational cost
Our USAC algorithm uses twin critics and a single actor network as for SAC, TD3, OAC, and TOP. This design choice eliminates the need for any additional computational steps and is no reliance on repetitive sampling, pre-training, or iterative approximation procedures. Consequently, it can be conclude that its computational characteristics closely match those of other widely used actor-critic algorithms and ours comparing baselines reported. The code for our algorithm implementation is also available alongside the paper.

### v2

@@
-Exploring Pessimism and Optimism Dynamics in Deep Reinforcement Learning
+Improving Actor-Critic Training with Steerable Action-Value Approximation Errors
 Abstract
-Off-policy actor-critic algorithms have shown promise in deep reinforcement learning for continuous control tasks.
-Their success largely stems from leveraging pessimistic state-action value function updates, which effectively address function approximation errors and improve performance.
-However, such pessimism can lead to under-exploration, constraining the agent’s ability to explore/refine its policies.
-Conversely, optimism can counteract under-exploration, but it also carries the risk of excessive risk-taking and poor convergence if not properly balanced.
-Based on these insights, we introduce Utility Soft Actor-Critic (USAC), a novel framework within the actor-critic paradigm that enables independent control over the degree of pessimism/optimism for both the actor and the critic via interpretable parameters.
-USAC adapts its exploration strategy based on the uncertainty of critics through a utility function that allows us to balance between pessimism and optimism separately.
-By going beyond binary choices of optimism and pessimism, USAC represents a significant step towards achieving balance within off-policy actor-critic algorithms.
-Our experiments across various continuous control problems show that the degree of pessimism or optimism depends on the nature of the task.
-Furthermore, we demonstrate that USAC can outperform state-of-the-art algorithms for appropriately configured pessimism/optimism parameters.
+Off-policy actor-critic algorithms have shown strong potential in deep reinforcement learning for continuous control tasks.
+Their success primarily comes from leveraging pessimistic state-action value function updates, which reduce function approximation errors and stabilize learning.
+However, excessive pessimism can limit exploration, preventing the agent from effectively refining its policies.
+Conversely, optimism can encourage exploration but may lead to high-risk behaviors and unstable learning if not carefully managed.
+To address this trade-off, we propose Utility Soft Actor-Critic (USAC), a novel framework that allows independent, interpretable control of pessimism and optimism for both the actor and the critic.
+USAC dynamically adapts its exploration strategy based on the uncertainty of critics using a utility function, enabling a task-specific balance between optimism and pessimism.
+This approach goes beyond binary choices of pessimism or optimism, making the method both theoretically meaningful and practically feasible.
+Experiments across a variety of continuous control tasks show that adjusting the degree of pessimism or optimism significantly impacts performance.
+When configured appropriately, USAC consistently outperforms state-of-the-art algorithms, demonstrating its practical utility and feasibility.
+[…]
+6112
 1 Introduction
-Deep reinforcement learning (RL) faces significant challenges when navigating the complex landscape of high-dimensional state spaces and non-linear state-action functions.
-The deadly triad of function approximation, off-policy learning, and bootstrapping compounds these challenges, leading to instability and poor sample efficiency [19, 27, 33].
-One particularly critical consequence is the overestimation bias [29], where the estimated state-action values ($Q$-values) exceed their true counterparts [31, 32].
-The overestimation bias is typically addressed through a pessimistic approach that use the minimum between two critics’ estimations [10, 12].
-While effective in reducing overestimation, this approach can lead to pessimistic under-exploration [7], which limits the agent’s ability to explore and discover new policies.
-A natural remedy for pessimistic under-exploration could involve integrating optimism into actors’ exploration; however, the balance between pessimism and optimism is not straightforward, as this balance varies across tasks and evolves during the training process [20].
-The core focus of this paper is to explore the dynamics between pessimism and optimism in deep RL, especially within actor critics.
-To address this dynamic, we introduce the novel framework termed Utility Soft Actor-Critic (USAC).
-Unlike binary approaches, USAC considers pessimism and optimism as points along a spectrum.
-Our aim is to address the overestimation bias while avoiding pessimistic under-exploration.
-In the next section, we explore the current literature on this topic and elaborate on our main contributions.
-2 Exploring pessimism and optimism in deep reinforcement learning
-The primary focus of this work is to understand the interplay between pessimism and optimism and their impact on overestimation bias and pessimistic under-exploration.
-In this section, we delve into related work addressing these challenges, exploring various strategies to overcome them.
-Moreover, we conclude this section by outlining our main contributions.
-Addressing overestimation bias with pessimism
-A pessimistic approach has proven effective in reducing the overestimation bias [^1] The overestimation bias stems from inherent noise and approximation errors introduced by the function approximations and is further amplified by bootstrapping [29, 31, 32].
-– as demonstrated in Twin Delayed DDPG [^2] The Deep Deterministic Policy Gradient (DDPG) algorithm was proposed by Lillicrap et al.
-[17].
-(TD3) [10] – it also introduces challenges related to pessimistic under-exploration [7].
-Soft Actor-Critic (SAC) [12] addresses this challenge by integrating an entropy term into the actor’s policy decision-making process, thus making the actor stochastic.
-By incorporating the entropy, SAC encourages exploration and diversifies the actor’s actions, aiming to alleviate the adverse effects of pessimistic under-exploration.
-However, despite its effectiveness, entropy regularization alone do not resolve the underlying issues.
-Exploring optimism to counter pessimistic under-exploration
-An intuitive approach to tackle pessimistic under-exploration is based on a general form of the optimism in the face of uncertainty principle.
+Deep reinforcement learning (RL) faces significant challenges when navigating the complex landscape of high-dimensional state spaces and non-linear state-action functions, especially in continuous control tasks.
+The deadly triad of function approximation, off-policy learning, and bootstrapping compounds these challenges, often leading to instability and poor sample efficiency [21, 30, 36].
+One particularly critical consequence is overestimation bias [32], where the estimated action-values ($Q$-values) exceed their true counterparts [34, 35].
+Overestimation bias in off-policy RL can arise from multiple sources.
+A well-studied source is the maximization operator in temporal-difference learning, which systematically drives estimates upward by preferring positive noises.
+It is typically addressed by min-clipping [12, 14], where the minimum of two critic networks is used, though this introduces a systematic underestimation bias [34].
+Other contributors include sampling error in batch temporal-difference learning [24] and reuse bias, which arises when the same data are repeatedly used for updates [39].
+Although the remedies reduce overestimation, they often swing too far toward pessimism, producing overly pessimistic policy evaluations discouraging exploration.
+This phenomenon, known as pessimistic under-exploration [9], limits effective learning in actor-critic methods, necessitating methods that can balance optimism and pessimism more systematically.
+However, the interaction between critic errors and resulting actor behavior follows complex dynamics in online deep RL [22], and remains poorly understood theoretically.
+We posit that the impact of the interplay between optimism and pessimism on the learning profiles of actor-critic pipelines can be explained by a simple and intuitive pair of hyperparameters.
+We introduce a new framework that puts the action-value predictions of a pair of critics into a utility-theoretic perspective, which considers optimism and pessimism as points along a spectrum.
+This framework is practical as the hyperparameters controlling optimism and pessimism are interpretable and can be tuned for different tasks.
+The commonplace min-clipping boils down to only one of infinitely many possibilities within this spectrum.
+The descriptive value of this spectrum can be of the interest to theoreticians.
+It can also be used by practitioners as a hyperparameter to improve performance via grid-search.
+2 Prior art of orchestrating optimism and pessimism
+The state-of-the-art suggests the following three strategies to counteract the under-exploratory effects of pessimistic policy evaluation:
+(i) Maximum-entropy policy search.
+Algorithms such as Soft Actor-Critic (SAC) [14] maintain exploration by incentivizing an entropy increase in the policy distribution.
+This incentive is added to the reward function in the Bellman target calculation and is accounted for also during actor training.
+In effect, this is a random exploration scheme that perturbs the action selection independently from an estimation of the accuracy of the action-value approximations.
+This disconnect manifests itself as the brittleness of the algorithm on the entropy coefficient [15].
+(ii) Optimistic exploration.
+Many theoretically-grounded approaches for model-driven exploration adopt the Optimism in the Face of Uncertainty (OFU) principle [1, 2].
 Ciosek et al.
-[7] introduced the Optimistic Actor-Critic (OAC) algorithm, which adopts an optimistic approximate upper confidence bound on the $Q$-value function to enhance policy exploration.
-However, it’s crucial to note that aggressive exploration, as encouraged by optimistic algorithms like OAC, can be risky, particularly in environments where the size of overestimation is not well-understood [20].
-Figure 1: Impact of pessimism/optimism levels on critic and actor.
-Top: Final returns.
-Bottom: Estimation error (i.e., true - estimation); positive values means underestimation, whereas negative values means overestimation.
-Balancing between pessimism and optimism
-The balance of pessimism and optimism is nuanced and it varies across different tasks and develops throughout the training process [20].
+[9] implemented this principle into a deep actor-critic pipeline with their seminal Optimistic Actor-Critic (OAC) algorithm.
+OAC adopts an optimistic approximate upper confidence bound on the $Q$-value function to enhance policy exploration while using a pessimistic lower-confidence bound for critic training.
+The algorithm is computationally expensive.
+Furthermore, the effectiveness of its exploration is limited by the approximation errors incurred during confidence set calculation.
+(iii) Dynamic exploration.
+The balance of optimism and pessimism is nuanced, and it varies across different tasks and develops throughout the training process [22].
 The Tactical Optimism and Pessimism (TOP) algorithm, introduced by Moskovitz et al.
-[20], addresses this variability by dynamically balancing pessimism and optimism using multi-armed bandits [6, 16].
-However, unlike algorithms such as OAC, which lean towards optimism, TOP learns to adjust the degree of pessimism or optimism based on the prevailing uncertainty.
-However, it uses the same level of pessimism/optimism for both critic and actor training, which limits its versatility.
-USAC and our main contributions
-Inspired by these insights, we see pessimism and optimism as points along a spectrum rather than binary opposites.
-While TOP is rooted in the TD3 algorithm, our approach stems from SAC.
-We introduce a novel framework called USAC, which quantifies the inherent uncertainty in learning through a utility function.
-Unlike binary approaches, USAC adapts its exploration strategy based on the critics’ distribution through uncertainty estimates of their $Q$-values.
-Moreover, USAC permits the fine-tuning of pessimism or optimism levels during critic or actor training.
-Through this methodology, USAC strives for a balanced pessimism-optimism trade-off that overcomes both overestimation bias and pessimistic under-exploration.
-Figure 1 illustrates the impact of varying levels of pessimism and optimism on both the critic and actor components.
-This indicates that optimal levels of pessimism and optimism can effectively address both overestimation (or underestimation) and pessimistic under-exploration, thereby enhancing the overall return.
-Through empirical evaluations and comparisons with state-of-the-art algorithms, we demonstrate the versatility and robustness of USAC across diverse continuous control tasks, paving the way for more adaptive and efficient learning algorithms in the field of deep RL.
-3 Preliminary
+[22], addresses this variability by dynamically balancing optimism and pessimism multi-armed bandits [8, 18].
+The approach assumes a Gaussian distributed action-value function and it tunes the degree of exploration by playing a bandit game on a predetermined grid of quantiles.
+The TOP algorithm uses the same level of pessimism/optimism for both critic and actor training in order to maintain the feasibility of the bandit game, which limits the versatility of the algorithm.
+Furthermore, as the algorithm is designed upfront for on-the-fly tuning, it does not permit an interpretable analysis of pessimism and optimism in environments of interest.
+We propose a simpler alternative: treating the critic pairs as samples from a parametric distribution over action-value functions.
+For example, when using a Laplace distribution as our main focus in the paper, embedding it within a utility function allows us to express the degree of critic pessimism and actor optimism on an interpretable scale between 1-1 and +1+1.
+We demonstrate the expressive and instrumental power of this framework within a SAC extension due to the wide adoption of the model family.
+We refer to the resulting algorithm as a Utility Soft Actor Critic (USAC).
+However, our framework is equally applicable to alternative families such as deterministic policy gradient variants [19, 12].
+As shown in Figure˜1, our framework reveals how the specific demands of an environment for pessimistic policy evaluation and optimistic policy search influence each other.
+The left panel illustrates the impact of different combinations of pessimism and optimism on final returns, where a coarse grid-search can yield improvements exceeding 30% in several cases.
+The right panel shows the corresponding effect on value estimation bias (true value minus estimated value), indicating whether the policy tends to under- or over-estimate returns.
+Figure 1: Impact of optimism and pessimism levels on actor–critic performance on Walker2d environment.
+Left: Final returns for different combinations of pessimism and optimism.
+Right: Value estimation bias (i.e., true – estimated value); positive values indicate underestimation, negative values indicate overestimation
+3 Preliminaries
 Throughout this paper, we denote $\mathcal{P}(\Omega)$ as the set of all probability distributions on $\Omega$ and let $\mathcal{B}(\Omega)$ be the set of bounded functions on $\Omega$.
-Markov decision processes (MDPs)
-[…]
-Policies
+Markov decision processes (MDPs).
+An MDP can be represented by the tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},p,p_{0},r,\gamma\rangle$ [26].
+Here, $\mathcal{S}$ and $\mathcal{A}$ denote the continuous state and action spaces.
+The function $p(s^{\prime}|s,a)$ for $(s,a,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}$ represents the transition probability from the current state $s$ and action $a$ to the next state $s^{\prime}$.
+We consider the model-free setup where this function is not known and will not be learned.
+The initial state distribution is denoted by $p_{0}\in\mathcal{P}(\mathcal{S})$, while $r:\mathcal{S}\times\mathcal{A}\rightarrow[0,B_{r}]$ represents the bounded reward function with $B_{r}>0$, and $\gamma\in[0,1]$ stands for the discount factor.
+Policies.
 Let $\Pi=\{\pi:\mathcal{S}\rightarrow\mathcal{P}(\mathcal{A})\}$ be the set of policies.
-The interaction between the agent and the MDP $\mathcal{M}$ under some policy $\pi\in\Pi$ progresses iteratively: At each time step $t\in\mathbb{N}$, the agent observes state $s_{t}\in\mathcal{S}$, chooses action $a_{t}\in\mathcal{A}$ based on the policy $a_{t}\sim\pi(\cdot|s_{t})$, receives a (bounded) reward $r_{t}:=r(s_{t},a_{t})$, and transitions to the next state $s_{t+1}\sim p(\cdot|s_{t},a_{t})$.
+The interaction between the agent and the MDP $\mathcal{M}$ under some policy $\pi\in\Pi$ progresses iteratively.
+At each time step $t\in\mathbb{N}$, the agent observes state $s_{t}\in\mathcal{S}$, chooses action $a_{t}\in\mathcal{A}$ based on the policy $a_{t}\sim\pi(\cdot|s_{t})$, receives a (bounded) reward $r_{t}:=r(s_{t},a_{t})$, and transitions to the next state $s_{t+1}\sim p(\cdot|s_{t},a_{t})$.
 Additionally, we define $p^{\pi}(s^{\prime},a^{\prime}|s,a)=p(s^{\prime}|s,a)\pi(a^{\prime}|s^{\prime})$ as the one-step transition probability from $(s,a)$ to $(s^{\prime},a^{\prime})$.
-Maximum entropy RL
+Maximum entropy RL.
 The standard goal of RL is to find a policy $\pi$ that maximizes the expected sum of discounted rewards $J(\pi)=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$ with initial state $s_{0}\sim p_{0}$ [30, 5, 31].
@@
 Formally, this is achieved by incorporating an entropy term, tempered by a parameter $\alpha>0$, which governs the relative importance of the return compared to the policy entropy.
-The maximum entropy objective function is defined as $J_{\alpha}(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}(r_{t}+\alpha\mathcal{H}(\pi(\cdot|s_{t})))\right]$, where the policy entropy term $\mathcal{H}(\pi(\cdot|s))=-\mathbb{E}_{a\sim\pi(\cdot|s)}[\log\pi(a|s)]$.
+The maximum entropy objective function is defined as $J_{\alpha}(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}(r_{t}+\alpha\mathcal{H}(\pi(\cdot|s_{t})))\right]$}, where the policy entropy term is defined as $\mathcal{H}(\pi(\cdot|s))=-\mathbb{E}_{a\sim\pi(\cdot|s)}[\log\pi(a|s)]$.
 […]
-$T^{\pi}Q(s,a)=r(s,a)+\gamma\mathbb{E}_{(s^{\prime},a^{\prime})\sim p^{\pi}(\cdot|s,a)}[Q(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})]\,,$ (1)
-i.e., $T^{\pi}Q^{\pi}(s,a)=Q^{\pi}(s,a)$ for any $(s,a)\in\mathcal{S}\times\mathcal{A}$.
-Actor-critic algorithms
-One approach to learning the optimal policy is through the framework of actor-critic algorithms.
-In these algorithms, an agent simultaneously learns a policy (the actor), responsible for selecting actions that maximize expected return, and an estimation of the state-action value function (the critic), responsible for evaluating the policy’s quality through iterative updates.
-Formally, the critic $Q$ estimates the state-action value function $Q^{\pi}$ (derived by following policy $\pi$) by
+[…]
+that is, $T^{\pi}Q^{\pi}(s,a)=Q^{\pi}(s,a)$ for any $(s,a)\in\mathcal{S}\times\mathcal{A}$.
+Actor-critic algorithms.
+Performing continuous control with deep RL necessitates a policy-gradient approach applicable to online updates.
+The actor-critic design is an optimal and widely adopted framework for this purpose.
+In these algorithms, an agent simultaneously learns a policy (the actor), responsible for selecting actions that maximize expected return, and an estimation of the action-value function (the critic), responsible for evaluating the policy’s quality through iterative updates.
+Formally, the critic $Q$ is trained to predict the action-value function $Q^{\pi}$ (derived by following policy $\pi$) by temporal difference learning
 $\operatorname*{arg\,min}_{Q}\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}[(Q(s,a)-y(s,a,r,s^{\prime}))^{2}],$ (2)
@@
 $\operatorname*{arg\,max}_{\pi\in\Pi}\mathbb{E}_{a\sim\pi(\cdot|s)}[\tilde{Q}(s,a)-\alpha\log\pi(a|s)],$ (4)
-where $\tilde{Q}$ is (again) an estimate of the state-action value function $Q^{\pi}$, which can differ from both the critic $Q$ and the target critic $\bar{Q}$.
-In scenarios where the agent do a full Bellman backup without approximation errors, the critic $Q$, the target critic $\bar{Q}$, and the critic used to update the actor $\tilde{Q}$ are all identical to $Q^{\pi}$ [27, 12].
-However, variations in $Q$, $\bar{Q}$, and $\tilde{Q}$ are often introduced to address the challenges we outlined in Section 2; we will explore some of these variations in more details below.
-Exploring pessimism and optimism in actor-critic algorithms
-A common strategy to address overestimation involves estimating the state-action value function using twin critics, along with associated target critics that are delayed versions of these critics [17, 18, 19, 10].
+where $\tilde{Q}$ is (again) an estimate of the action-value function $Q^{\pi}$, which can differ from both the critic $Q$ and the target critic $\bar{Q}$.
+In scenarios where the agent does a full Bellman backup without approximation errors, the critic $Q$, the target critic $\bar{Q}$, and the critic used to update the actor $\tilde{Q}$ are all identical to $Q^{\pi}$ [30, 14].
+However, variations in $Q$, $\bar{Q}$, and $\tilde{Q}$ are often introduced to address the challenges outlined in the introduction.
+We will explore some of these variations in more detail below.
+The sources of optimism and pessimism in actor-critics.
+A common strategy to address overestimation involves estimating the action-value function using twin critics, along with associated target critics that are delayed versions of these critics [19, 20, 21, 12].
 Specifically, let $Q_{1},Q_{2}:\mathcal{S}\times\mathcal{A}\rightarrow{\mathbb{R}}$ be the two critics and $\bar{Q}_{1},\bar{Q}_{2}:\mathcal{S}\times\mathcal{A}\rightarrow{\mathbb{R}}$ the two target critics that are delayed versions of the critics.
@@
 Similarly, the same approach is employed to update the actor (in TD3 and SAC).
-Here, the critic to update the actor, $\tilde{Q}$ in (4), is determined by the minimum of the two critics $\{Q_{k}\}_{k=1,2}$:
+Here, the critic used to update the actor, $\tilde{Q}$ in (4), is determined by the minimum of the two critics $\{Q_{k}\}_{k=1,2}$:
 $\tilde{Q}(s,a)=\min_{k=1,2}Q_{k}(s,a).$ (6)
 Ciosek et al.
-[7] noted that the pessimistic approach, described in 5 and 6, can lead to pessimistic under-exploration.
+[9] noted that the pessimistic approach, described in ˜5 and 6, can lead to pessimistic under-exploration.
 To prevent this, they proposed embracing the optimism in the face of uncertainty principle in their OAC algorithm.
 However, this approach can potentially make exploration too aggressive, thereby negatively impacting overall performance [22].
-Instead, the balance between pessimism and optimism should be tailored to the problem’s nature.
+Instead, the balance between optimism and pessimism should be tailored to the problem’s nature.
 Moskovitz et al.
@@
 $\bar{Q}(s,a)=\bar{\mu}(s,a)+\beta\bar{\sigma}(s,a),\quad\beta\in{\mathbb{R}},$ (7)
-where the mean $\bar{\mu}$ and (unbiased) standard deviation $\bar{\sigma}$ are constructed using the two target critics $\bar{Q}_{1},\bar{Q}_{2}$; $\bar{\mu}(s,a)=\frac{1}{2}(\bar{Q}_{1}(s,a)+\bar{Q}_{2}(s,a))$ and $\bar{\sigma}(s,a)=\sqrt{\sum_{k=1,2}(\bar{Q}_{k}(s,a)-\bar{\mu}(s,a))^{2}}$.
+where the mean $\bar{\mu}$ and (unbiased) standard deviation $\bar{\sigma}$ are constructed using the two target critics $\bar{Q}_{1},\bar{Q}_{2}$:
+$\bar{\mu}(s,a)=\frac{1}{2}(\bar{Q}_{1}(s,a)+\bar{Q}_{2}(s,a))$
+and
+$\bar{\sigma}(s,a)=\sqrt{\sum_{k=1,2}(\bar{Q}_{k}(s,a)-\bar{\mu}(s,a))^{2}}.$
 The critic to update the actor, $\tilde{Q}$, can be defined in a similar manner from the two critics $\{Q_{k}\}_{k=1,2}$.
 Moskovitz et al.
-[20] classify 7 as optimistic when $\beta\geq 0$ and pessimistic when $\beta<0$.
-In the special case when $\beta=-1/\sqrt{2}$, 7 simplifies to taking the minimum of the two critics – as in 5 and 6.
-The TOP algorithm tries to adjust the balance between pessimism and optimism by using a multi-armed bandit algorithm to select $\beta$ from the set $\{-1,0\}$; here $\beta=-1$ corresponds to a pessimistic estimate and $\beta=0$ to an optimistic one.
+[22] classify ˜7 as optimistic when $\beta\geq 0$ and pessimistic when $\beta<0$.
+In the special case when $\beta=-1/\sqrt{2}$, ˜7 simplifies to taking the minimum of the two critics as in ˜5 and 6.
+The TOP algorithm tries to adjust the balance between optimism and pessimism by using a multi-armed bandit algorithm to select $\beta$ from the set $\{-1,0\}$; here $\beta=-1$ corresponds to a pessimistic estimate and $\beta=0$ to an optimistic one.
 Although $\beta$ can vary over time using a bandit algorithm, TOP uses the same $\beta$ for both $\bar{Q}$ and $\tilde{Q}$.
-4 Utility Soft Actor-Critic (USAC): framework and algorithms
-Unlike the aforementioned algorithms, our framework decouples the levels of optimism and pessimism for the actor and critic.
-This novel approach allows us to be pessimistic in learning the critic while being optimistic in actor training.
-By doing so, we address the overestimation bias observed during critic training and overcome the problem of pessimistic under-exploration faced by the actor.
-Consequently, our approach promotes both a stable critic and an explorative actor simultaneously.
-In this section, we establish the theoretical groundwork for exploring the balance between optimism and pessimism.
-We introduce the Utility Soft Actor-Critic (USAC) framework, which seamlessly integrates the uncertainty of critics into the actor learning process.
-Our framework guarantees policy improvement, as detailed in Section 4.1, while offering flexibility in navigating the optimism-pessimism trade-off.
-Moving on to Sections 4.2 and 4.3, we present the USAC algorithm and demonstrate how specific choices within our framework can emulate the optimism and pessimism strategies of algorithms like TD3 [10], SAC [12], OAC [7], and TOP [20]; in Section 5, we will explore these varying levels of optimism and pessimism across five MuJoCo environments.
-The pseudo-code of USAC is provided in Algorithm 1.
-Before opening the entire discussion, we first introduce how to quantify the uncertainty of the $Q$-functions.
-Quantifying the uncertainty of the $Q$-value distribution
+4 A utility-theoretic framework to make the action-value biases steerable
+Our key contribution is a new theoretical framework to quantify the degree of optimism or pessimism assumed by an actor-critic algorithm that performs policy evaluation via twin critics using two interpretable scalars, one for the critic and one for the actor.
+We achieve this by assuming that the action-value function follows a parametric distribution and the twin critics are two samples from this distribution.
+The resulting utility function yields a form that enables describing the optimism and pessimism levels via scalar risk coefficients that take the commonplace min-clipping method as a special case and provides an effortless generalization of it by simple grid-search.
 Let $\mathcal{Q}\in\mathcal{P}(\mathcal{B}(\mathcal{S}\times\mathcal{A}))$ be a distribution over $Q$-functions, serving as the agent’s estimate of the $Q$-functions.
@@
 This utility function is commonly used for measuring risk in finance, economics, and decision-making under uncertainty [11, 37].
-[^3] The utility function in 8 is also referred to as expected utility, exponential utility, exponential risk measure, generalized mean, or entropic risk measure according to the context [8, 9].
 It also shares similarities with the utility functions employed in risk-sensitive RL [16, 29, 25] and distributional RL [27, 3, 4].
-It’s important to note that we apply this utility function to the $\mathcal{Q}$ distribution rather than the return distribution, as is typically done in risk-sensitive and distributional RL.
-In this work, we use 8 as a measure of pessimism or optimism of the $\mathcal{Q}$ distribution: a positive value of $\lambda$ implies an optimistic view, while a negative value indicates a pessimistic view.
-This interpretation is clear from the Taylor expansion: $U^{\mathcal{Q}}_{\lambda}=\mathbb{E}_{Q\sim\mathcal{Q}}[Q]+\frac{\lambda}{2}\mathbb{V}_{Q\sim\mathcal{Q}}[Q]+\mathcal{O}(\lambda^{2})$.
+It is noteworthy that we apply this utility function to the $\mathcal{Q}$ distribution rather than the return distribution, as is typically done in risk-sensitive and distributional RL.
+The utility function is also referred to as expected utility, exponential utility, exponential risk measure, generalized mean, or entropic risk measure according to the context [10, 11].
+We use ˜8 as a measure of pessimism or optimism of the $\mathcal{Q}$ distribution: a positive value of $\lambda$ implies an optimistic view, while a negative value indicates a pessimistic view.
+This interpretation is clear from a second-order Taylor expansion applied to the utility function:
+$U^{\mathcal{Q}}_{\lambda}=\mathbb{E}_{Q\sim\mathcal{Q}}[Q]+\frac{\lambda}{2}\mathbb{V}_{Q\sim\mathcal{Q}}[Q]+\mathcal{O}(\lambda^{2}).$
 Consequently, the utility function $U^{\mathcal{Q}}_{\lambda}$ spans the entire spectrum of the $\mathcal{Q}$ distribution by varying $\lambda$.
-Another important property of the utility is that when $\mathcal{Q}$ is a Dirac delta distribution centered at $Q^{\pi}$, the $Q$-function of a given policy $\pi\in\Pi$, the expected utility function $U^{\mathcal{Q}}_{\lambda}$ is equivalent to $Q^{\pi}$ for any finite $\lambda$.
-4.1 USAC framework: policy evaluation, improvement and guarantees
-In this section, we will explain the policy evaluation and policy improvement steps, along with the corresponding guarantees for our USAC framework.
-[…]
-$T^{\pi}U^{\mathcal{Q}}_{\lambda}(s,a)-U^{\mathcal{Q}}_{\lambda}(s,a)=r(s,a)+\gamma\mathbb{E}_{(s^{\prime},a^{\prime})\sim p^{\pi}(\cdot|s,a)}\left[U_{\lambda}^{\mathcal{Q}}(s^{\prime},a^{\prime})-\alpha\log(\pi(a^{\prime}|s^{\prime}))\right]-U_{\lambda}^{\mathcal{Q}}(s,a)$
-should be diminished. Nevertheless, it fosters exploration during learning by maintaining a distribution over the $Q$-functions, which can guide exploration.
-On the other hand, given a $\mathcal{Q}$ distribution associated with the current policy $\pi$, denoted as $\mathcal{Q}^{\pi}$, we update the policy, similar to 4, through the following maximization problem:
+4.1 Policy evaluation, improvement and guarantees
+For any given policy $\pi\in\Pi$, the utility $U^{\mathcal{Q}}_{\lambda}\in\mathcal{B}(\mathcal{S}\times\mathcal{A})$ will reduce to $Q^{\pi}$ for any $|\lambda|<\infty$, when $\mathcal{Q}$ is a Dirac delta distribution centered at $Q^{\pi}$.
+This implies that the deviation of $U^{\mathcal{Q}}_{\lambda}$ from $Q^{\pi}$ is due to the uncertainty in the learned $Q$-function.
+In fact, we observe that the Bellman operator should also bring the utility $U^{\mathcal{Q}}_{\lambda}$ to its fixed point $Q^{\pi}$ [5]
+Lemma 1.
+For any expected utility of $\mathcal{Q}\in\mathcal{P}(\mathcal{S}\times\mathcal{A})$, we have
+$\lVert T^{\pi}U^{\mathcal{Q}}_{\lambda}-U^{\mathcal{Q}}_{\lambda}\rVert^{2}_{\rho^{\pi}}\leq\mathbb{E}_{(s,a)\sim\rho^{\pi}}\mathbb{E}_{(s^{\prime},a^{\prime})\sim p^{\pi}(\cdot|s,a)}\left[\ell(U^{\mathcal{Q}}_{\lambda};\xi)\right],$
+where
+$\ell(U^{\mathcal{Q}}_{\lambda};\xi)=(r(s,a)+\gamma(U^{\mathcal{Q}}_{\lambda}(s^{\prime},a^{\prime})-\alpha\log(\pi(a^{\prime}|s^{\prime})))-U^{\mathcal{Q}}_{\lambda}(s,a))^{2}$
+is the squared loss of $U^{\mathcal{Q}}_{\lambda}$ on an observation $\xi=(s,a,s^{\prime},a^{\prime})$.
+Proof.
+For any $\mathcal{Q}\in\mathcal{P}(\mathcal{S}\times\mathcal{A})$, we have
+$\lVert T^{\pi}U^{\mathcal{Q}}_{\lambda}-U^{\mathcal{Q}}_{\lambda}\rVert_{\rho^{\pi}}^{2}=\mathbb{E}_{(s,a)\sim\rho^{\pi}}[(T^{\pi}U^{\mathcal{Q}}_{\lambda}(s,a)-U^{\mathcal{Q}}_{\lambda}(s,a))^{2}]$
+$=\mathbb{E}_{(s,a)\sim\rho^{\pi}}[(r(s,a)+\gamma\mathbb{E}_{(s^{\prime},a^{\prime})\sim p^{\pi}(\cdot|s,a)}[U^{\mathcal{Q}}_{\lambda}(s^{\prime},a^{\prime})$-
+$\quad\quad\quad\quad\quad\quad\alpha\log(\pi(a^{\prime}|s^{\prime}))]-U^{\mathcal{Q}}_{\lambda}(s,a))^{2}]$
+$\leq\mathbb{E}_{(s,a)\sim\rho^{\pi}}\mathbb{E}_{(s^{\prime},a^{\prime})\sim p^{\pi}(\cdot|s,a)}[(r(s,a)+\gamma(U^{\mathcal{Q}}_{\lambda}(s^{\prime},a^{\prime})$-
+$\quad\quad\quad\quad\quad\quad\alpha\log(\pi(a^{\prime}|s^{\prime})))-U^{\mathcal{Q}}_{\lambda}(s,a))^{2}],$
+where the inequality comes from the Jensen’s inequality. ∎
+Thus, as a consequence of Lemma˜1, the Bellman error defined by
+$T^{\pi}U^{\mathcal{Q}}_{\lambda}(s,a)-U^{\mathcal{Q}}_{\lambda}(s,a)=r(s,a)$
+$\qquad+\gamma\mathbb{E}_{(s^{\prime},a^{\prime})\sim p^{\pi}(\cdot|s,a)}\left[U_{\lambda}^{\mathcal{Q}}(s^{\prime},a^{\prime})-\alpha\log(\pi(a^{\prime}|s^{\prime}))\right]-U_{\lambda}^{\mathcal{Q}}(s,a)$
+should diminish.
+Nevertheless, it fosters exploration during learning by maintaining a distribution over the $Q$-functions.
+On the other hand, given a $\mathcal{Q}$ distribution associated with the current policy $\pi$, denoted as $\mathcal{Q}^{\pi}$, we update the policy, similar to ˜4, through the following maximization problem:
 $\operatorname*{arg\,max}_{\pi\in\Pi}\mathbb{E}_{a\sim\pi(\cdot|s)}[U_{\lambda}^{\mathcal{Q}^{\pi}}(s,a)-\alpha\log\pi(a|s)].$
 Extending the policy iteration process from SAC to utility functions ensures that the guarantees of policy improvement still hold [14, 13].
-4.2 USAC algorithm
-The USAC learning process follows the general actor-critic algorithm.
-In particular, we estimate the desired quantities by samples from a replay buffer $(s,a,r,s^{\prime})\sim\mathcal{D}$.
-As the operator $T^{\pi}$ should bring the utility $U^{\mathcal{Q}}_{\lambda}$ to its fixed-point $Q^{\pi}$, the USAC framework learns the critic $Q$ (an estimate of $Q^{\pi}$) as in 2 but with the critic’s target value defined by
-$y(s,a,r,s^{\prime})=r+\gamma[U_{\lambda_{\text{critic}}}^{\bar{\mathcal{Q}}}(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})],\quad a^{\prime}\sim\pi(\cdot|s^{\prime}),$ (9)
-which is built using a target critic distribution $\bar{\mathcal{Q}}$ with some utility parameter $\lambda_{\text{critic}}$ chosen specifically for the critic.
-Similar to 4, the actor learns its policy by
+4.2 The Utility Soft Actor Critic (USAC)
+We illustrate a simple implementation of our utility-theoretic formulation into a soft actor-critic pipeline and refer to the resulting algorithm as the Utility Soft Actor Critic (USAC).
+We perform online and off-policy learning as usual, maintaining a replay buffer of observed environment interactions $(s,a,r,s^{\prime})\sim\mathcal{D}$.
+As the operator $T^{\pi}$ should bring the utility $U^{\mathcal{Q}}_{\lambda}$ to its fixed-point $Q^{\pi}$, the USAC framework learns the critic $Q$ (an estimate of $Q^{\pi}$) as in ˜2 but with the critic’s target value defined as
+$y(s,a,r,s^{\prime})=r+\gamma[U_{\lambda_{\text{critic}}}^{\bar{\mathcal{Q}}}(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})],$ (9)
+where $a^{\prime}\sim\pi(\cdot|s^{\prime})$.
+Note that we use a target critic distribution $\bar{\mathcal{Q}}$ with some utility parameter $\lambda_{\text{critic}}$ chosen specifically for the critic.
+Similar to ˜4, the actor learns its policy by
 $\operatorname*{arg\,max}_{\pi\in\Pi}\mathbb{E}_{a\sim\pi(\cdot|s)}[U_{\lambda_{\text{actor}}}^{\tilde{\mathcal{Q}}}(s,a)-\alpha\log\pi(a|s)],$ (10)
 using $\tilde{\mathcal{Q}}$ as the estimate for $\mathcal{Q}$ and $\lambda_{\text{actor}}$ as the utility parameter for the actor.
-Note that although Moskovitz et al.
-[20] consider an optimism/pessimism parameter that can switch between two values over time, they use the same parameter for both actor and critic training.
-In contrast, our work decouples the optimism/pessimism levels of the actor and critic.
-This novel approach allows us to select a pessimistic parameter $\lambda_{\text{critic}}$ for learning the critic while using an optimistic parameter $\lambda_{\text{actor}}$ for training the actor.
-This strategy effectively addresses the overestimation bias observed during critic training and mitigates the issue of pessimistic under-exploration encountered by the actor.
-Consequently, our approach promotes a stable critic and an explorative actor simultaneously.
-Later in the experiments presented in Section 5, we will explore various combinations of pessimism and optimism levels.
-Before proceeding, we will ground the algorithm by considering a specific distribution.
-It’s worth noting that while we focus on the Laplace distribution in this paper, the framework described in this section is applicable to other distributions as well.
-The specific choice of Laplace is simply because it makes it more tangible to fine-tune the level of pessimism/optimism in the critic and actor, and to compare with other actor-critic algorithms.
-4.3 USAC algorithm: Laplace distribution
-The utility formulation in 8 enables us to compute the utility for any distribution with a moment-generating function.
-Let’s consider a specific case where $\mathcal{Q}$ follows a Laplace distribution.
-The algorithm overview is provided in Algorithm 1.
+Note that Moskovitz et al.
+[22] implements a single shared optimism/pessimism parameter for the actor and the critic.
+In contrast, our framework naturally decouples the optimism and pessimism levels of the actor and critic into two separate scalars, $\lambda_{\text{critic}}$ and $\lambda_{\text{actor}}$, respectively.
+This decoupling allows both investigating the relationship between the optimism and pessimism demands of the actor and the critic and tuning them separately, thereby benefiting from their complementary roles in online training.
+Later in the experiments presented in Section˜5, we will explore various combinations of optimism and pessimism levels.
+Algorithm 1 USAC with Laplace distribution (Section˜4.3)
+Averaging parameter $\tau\in(0,1)$, learning rates $\eta_{\theta},\eta_{\phi}>0$, mini-batch size $n\in\mathbb{N}$, pessimism/optimism coefficients $\kappa_{\text{actor}},\kappa_{\text{critic}}\in(-1,1)$, entropy coefficient $\alpha$,
+1:Polyak coefficient $\tau$
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+4.3 Laplace distributed critics
+The utility formulation in ˜8 enables us to compute the utility for any distribution with a moment-generating function.
+We will next present an implementation of a specific case of our USAC algorithm where $\mathcal{Q}$ follows a Laplace distribution, which has the desirable property of expressing the utility parameter within the [1,+1][-1,+1].
+The pseudo-code of the resulting USAC algorithm is provided in Algorithm˜1.
 Here, we denote the utility parameter associated with this distribution as $\kappa$, reflecting its characteristics within the Laplace context.
@@
 where $\mu_{\mathcal{Q}}$ and $\sigma_{\mathcal{Q}}$ represent the mean and standard deviation of $\mathcal{Q}$, respectively, and
-$g(\kappa)=\begin{cases}\log(1/(1-\kappa^{2}))/\sqrt{2}\kappa&\>\text{for}\>\kappa\in(-1,1)\backslash\{0\},\\ 0&\>\text{for}\>\kappa=0.\end{cases}$ (12)
-The proof of Proposition 1 can be found in Appendix A.
-Additionally, we present a corresponding proposition for Gaussian-distributed $Q$-values in Appendix B.
-In Proposition 1, we observe that by varying $\kappa$ in the small interval $(-1,1)$, $U^{\mathcal{Q}}_{\kappa}$ spans the entire spectrum of the $\mathcal{Q}$ distribution.
-Specifically, as $\lim_{\kappa\rightarrow-1}g(\kappa)\rightarrow-\infty$, $\lim_{\kappa\rightarrow 0}g(\kappa)\rightarrow 0$, and $\lim_{\kappa\rightarrow 1}g(\kappa)\rightarrow\infty$.
-Now suppose the uncertainty follows a Laplace distribution.
-Following the algorithmic steps outlined in Section 4.2, we can estimate the utility of the target critic distribution $\bar{\mathcal{Q}}$ by two target critics $\{\bar{Q}_{k}\}_{k=1,2}$, which are delayed versions of the critics $\{Q_{k}\}_{k=1,2}$.
-The utility function is given by $U^{\bar{\mathcal{Q}}}_{\kappa_{\text{critic}}}=\mu_{\bar{\mathcal{Q}}}+g(\kappa_{\text{critic}})\sigma_{\bar{\mathcal{Q}}}$ with utility parameter $\kappa_{\text{critic}}\in(-1,1)$, where
-$\mu_{\bar{\mathcal{Q}}}(s,a)=\frac{1}{2}(\bar{Q}_{1}(s,a)+\bar{Q}_{2}(s,a))\quad\text{and}\quad\sigma_{\bar{\mathcal{Q}}}(s,a)=\frac{1}{2}\lvert\bar{Q}_{1}(s,a)-\bar{Q}_{2}(s,a)\rvert.$ (13)
-Hence, the target value to train the critic becomes
-$y(s,a,r,s^{\prime})=r+\gamma[\mu_{\bar{\mathcal{Q}}}(s^{\prime},a^{\prime})+g(\kappa_{\text{critic}})\sigma_{\bar{\mathcal{Q}}}(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})],\quad a^{\prime}\sim\pi(\cdot|s^{\prime}).$ (14)
-Similarly, the actor learns its policy by
-$\operatorname*{arg\,max}_{\pi\in\Pi}\mathbb{E}_{a\sim\pi(\cdot|s)}[\mu_{\tilde{\mathcal{Q}}}(s,a)+g(\kappa_{\text{actor}})\sigma_{\tilde{\mathcal{Q}}}(s,a)-\alpha\log\pi(a|s)],\quad\kappa_{\text{actor}}\in(-1,1),$ (15)
-where $\mu_{\tilde{\mathcal{Q}}}$ and $\sigma_{\tilde{\mathcal{Q}}}$ are estimated in similar manner as 13 on $\tilde{\mathcal{Q}}$ using the critics $\{Q_{k}\}_{k=1,2}$.
-The pseudo-code of this algorithm can be found in Algorithm 1.
-We note that a smaller $\kappa$ value indicates more pessimism, while a larger $\kappa$ value indicates more optimism.
-By choosing $\kappa$ such that $g(\kappa)=-1$ (i.e., $\kappa\approx-0.83$), the utility in 11 simply reduces to the minimum of the two critics.
-Thus, choosing $\kappa_{\text{critic}}=\kappa_{\text{actor}}\approx-0.83$, we obtain the pessimism choices of SAC [12] and TD3 [10].
-On the other hand, by choosing $\kappa_{\text{critic}}\approx-0.83$ and a positive $\kappa_{\text{actor}}$, we have a pessimism critic and an optimistic actor as explored in OAC [7].
-Lastly, by letting a multi-armed bandit algorithm choose $\kappa_{\text{critic}}=\kappa_{\text{actor}}\in\{-0.916563,0\}$, we recover the pessimistic/optimistic design suggested by TOP [20].
-[^4] To recover the pessimistic case of the TOP algorithm, we can choose $\kappa=-0.916563$, which corresponds to having the function $g$ in 12 equal to $\sqrt{2}$.
-We note that by choosing $-0.916563$, TOP considers a more pessimistic critic target than SAC, TD3, and OAC.
-In Section 5, we will explore various $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ pairs and discuss whether the previous wisdom about the effectiveness of these strategies holds across different environments.
-Algorithm 1 USAC with Laplace distribution (Section 4.3)
-1:Input: Averaging parameter $\tau\in(0,1)$, learning rates $\eta_{\theta},\eta_{\phi}>0$, mini-batch size $n\in\mathbb{N}$, pessimism/optimism coefficients $\kappa_{\text{actor}},\kappa_{\text{critic}}\in(-1,1)$
-2:Initialize: Replay buffer $\mathcal{D}=\varnothing$, initial state $s_{0}\sim p_{0}$, critic $\{\theta_{k}\}_{k=1,2}$, target critic $\{\bar{\theta}_{k}\}_{k=1,2}$ and actor policy $\phi$
-3:for each time step do
-4:     for each environment step do
-5:         $a_{t}\sim\pi_{\phi}(\cdot|s_{t})$triangleright Sample action from the policy $\phi$
-6:         $s_{t+1}\sim p(\cdot|s_{t},a_{t})$triangleright Sample transition from the environment
-7:         $\mathcal{D}\leftarrow\mathcal{D}\cup(s_{t},a_{t},r_{t},s_{t+1})$triangleright Store transition in the replay pool
-8:     end for
-9:     for each training step do
-10:         $\{(s_{i},a_{i},r_{i},s_{i}^{\prime},a_{i}^{\prime}):(s_{i},a_{i},r_{i},s_{i}^{\prime})\sim\mathcal{D},a_{i}^{\prime}\sim\pi_{\phi}(\cdot|s_{i}^{\prime})\}_{i=1}^{n}$triangleright Sample mini-batch
-11:         for $i\in\{1,\dots,n\}$ do $\triangleright$ Compute target critic distribution $\bar{\mathcal{Q}}$
-12:              $\mu_{\bar{\mathcal{Q}}}(s_{i},a_{i})\leftarrow\frac{1}{2}(Q_{\bar{\theta}_{1}}(s_{i},a_{i})+Q_{\bar{\theta}_{2}}(s_{i},a_{i}))$
-13:              $\sigma_{\bar{\mathcal{Q}}}(s_{i},a_{i})\leftarrow\frac{1}{2}\lvert Q_{\bar{\theta}_{1}}(s_{i},a_{i})-Q_{\bar{\theta}_{2}}(s_{i},a_{i})\rvert$
-14:         end for
-15:         for each critic, $k\in\{1,2\}$ do
-16:              $\theta_{k}\leftarrow\theta_{k}-\eta_{\theta}\nabla_{\theta_{k}}\left\{\frac{1}{n}\sum_{i=1}^{n}\left(Q_{\theta_{k}}(s_{i},a_{i})-y(s_{i},a_{i},r_{i},s_{i}^{\prime})\right)^{2}\right\}$triangleright Update critic
-17:             with $y(s_{i},a_{i},r_{i},s_{i}^{\prime})=r_{i}+\gamma[\mu_{\bar{\mathcal{Q}}}(s_{i}^{\prime},a_{i}^{\prime})+g(\kappa_{\text{critic}})\sigma_{\bar{\mathcal{Q}}}(s_{i}^{\prime},a_{i}^{\prime})-\alpha\log\pi_{\phi}(a_{i}^{\prime}|s_{i}^{\prime})]$
-18:         end for
-19:         for $i\in\{1,\dots,n\}$ do $\triangleright$ Compute actor distribution $\tilde{\mathcal{Q}}$
-20:              $\mu_{\tilde{\mathcal{Q}}}(s_{i},a_{i})\leftarrow\frac{1}{2}(Q_{\theta_{1}}(s_{i},a_{i})+Q_{\theta_{2}}(s_{i},a_{i}))$
-21:              $\sigma_{\tilde{\mathcal{Q}}}(s_{i},a_{i})\leftarrow\frac{1}{2}\lvert Q_{\theta_{1}}(s_{i},a_{i})-Q_{\theta_{2}}(s_{i},a_{i})\rvert$
-22:         end $\triangleright$ Update actor
-23:         $\phi\leftarrow\phi+\eta_{\phi}\nabla_{\phi}\left\{\frac{1}{n}\sum_{i=1}^{n}\left(\mu_{\tilde{\mathcal{Q}}}(s_{i},a_{i})+g(\kappa_{\text{actor}})\sigma_{\tilde{\mathcal{Q}}}(s_{i},a_{i})-\alpha\log\pi_{\phi}(a_{i}|s_{i})\right)\right\}$
-24:         for each critic, $k\in\{1,2\}$ do
-25:              $\bar{\theta}_{k}\leftarrow\tau\theta_{k}+(1-\tau)\bar{\theta}_{k}$triangleright Update target critic
-26:         end for
-27:     end for
-28:end for
-5 Experiments
-The goal of our experiments is to explore the interplay between critic and actor pessimism/optimism, estimation error (i.e., overestimation/underestimation), and performance.
-We compare our USAC algorithm with Laplace distribution (Section 4.3) with prior off-policy actor-critic algorithms.
-Specifically, we consider SAC [12], TD3 [10], OAC [7] and TOP [20] as baselines, as they represent the most relevant algorithms that incorporate optimism and pessimism.
-For SAC and USAC, we use auto-tuned entropy temperature $\alpha$ [13], while for OAC and TOP, we adopt the best parameters reported in their respective papers.
-We conduct our experiments on 5 continuous control environments in the MuJoCo physics engine [30].
-Each experiment is repeated across 5 different seeds, each comprising one million time steps, and evaluated over 10 episodes.
-The reported results represent the average (along with the standard deviation) across these repetitions.
-Additional details about architectures, baselines, computational cost, environment properties, grid search, hyper-parameters, learning curves, area under learning curves, estimation error plots and their values can be found in Appendix C.
-Insights on USAC performance
-In Table 1, we present two sets of results for our USAC algorithm: a default and a best pair of parameters $(\kappa_{\text{critic}},\kappa_{\text{actor}})$; we will discuss how these parameters are selected below.
-These results demonstrate that our USAC algorithm can outperform (or match) all baseline algorithms across all environments.
-This success reaffirms our hypothesis, as outlined in Section 4, that the levels of pessimism and optimism should vary between the critic and actor, as well as across environments [20].
-In particular, our findings reveals that there exist pairs of utility parameters, $(\kappa_{\text{critic}},\kappa_{\text{actor}})$, capable of overcoming overestimation while facilitating exploration, thus enhancing overall performance.
-In addition, USAC smaller standard deviation indicates an improved model stability during learning.
-The corresponding learning curves and estimation error curves for these results are presented in Figures 5 and 6 (Appendix C).For the Humanoid-v4 environment, we found that the default parameter pair outperformed the best pair identified through grid search, e.g., see Figure 2.
-Therefore, we used the default pair for this environment in our results (Table 1).
-However, we conjecture that by expanding our (limited) grid search, a better pair may exist.
-Table 1: Final return on MuJoCo environments trained with 1M time steps, averaged over 5 seeds.
-The best algorithms are highlighted in bold.
-$\pm$ corresponds to the standard deviation across repetitions.
-The default $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ are both $-0.831559$.
-The best $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ are listed in Table 2.
-Environment USAC (ours) SAC TD3 OAC TOP
-Default $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ Best $(\kappa_{\text{critic}},\kappa_{\text{actor}})$
-Ant-v4 $\boldsymbol{5139\pm 978}$ $\boldsymbol{5158\pm 1186}$ $4756\pm 1411$ $4091\pm 1303$ $4177\pm 1392$ $4334\pm 1276$
-HalfCheetah-v4 $\boldsymbol{11024\pm 849}$ $\boldsymbol{11736\pm 317}$ $10763\pm 895$ $10570\pm 1400$ $8684\pm 1678$ $7311\pm 3074$
-Hopper-v4 $3194\pm 810$ $\boldsymbol{3442\pm 247}$ $3185\pm 537$ $1986\pm 1154$ $3293\pm 113$ $\boldsymbol{3367\pm 154}$
-Humanoid-v4 $\boldsymbol{5602\pm 505}$ $5602\pm 505^{\star}$ $\boldsymbol{5503\pm 373}$ $5149\pm 737$ $5390\pm 234$ $5332\pm 445$
-Walker2d-v4 $\boldsymbol{4525\pm 534}$ $\boldsymbol{4530\pm 767}$ $3757\pm 1282$ $4369\pm 601$ $3467\pm 1200$ $4317\pm 631$
-Exploring the dynamics of pessimistic and optimistic with USAC
-To explore the dynamics of pessimism and optimism in USAC, we conducted a grid search across the 5 continuous control environments in MuJoCo.
-More details about this grid search can be found in Appendix C.
-We examined critics ranging from pessimistic to semi-optimistic and actors from pessimistic to optimistic.
-Specifically, we used $\kappa_{\text{critic}}\in\{-0.831559,-0.5,-0.33\}$ [^5] Recall that $\kappa=-0.831559$ corresponds to $g(\kappa)=1$, simplifying the utility function in 11 to the minimum of the two critics.
-Therefore, by setting $\kappa_{\text{critic}}\approx-0.83$, we align with the pessimistic choices of SAC [12], TD3 [10], and OAC [7]., and $\kappa_{\text{actor}}\in\{-0.99,-0.5,0,0.5,0.99\}$.
-The rationale for choosing more pessimistic (to semi-optimistic) $\kappa_{\text{critic}}$ values is to counteract the overestimation bias.
-Given that USAC stems from the framework of SAC, SAC was included in our figures for comparison.
-Table 2: Best $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$ parameters for USAC.
-Environment $\kappa_{\text{critic}}$ $\kappa_{\text{actor}}$
-Ant-v4 $-0.831559$ $-0.99$
-HalfCheetah-v4 $-0.33$ $-0.50$
-Hopper-v4 $-0.831559$ $0.50$
-Humanoid-v4 $-0.831559$ $-0.831559^{\star}$
-Walker2d-v4 $-0.831559$ $0.0$
-By comparing the final returns (Figure 2) with the estimation error results (Figure 3) for the different grid combinations, we reaffirm previous findings [10, 7, 20]: i) overestimation (negative estimation error) should always be avoided as it reduces performance, and ii) underestimation (positive estimation error) does not necessarily harm performance.
-This indicates that a pessimistic approach should be adopted in critic training, whereas in actor training, either a pessimistic or optimistic approach can be suitable depending on the environment.
-The best combinations of $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$, with respect to final returns, are summarized in Table 2.
-Additionally, the area under the learning curve for the grid search parameters is illustrated in Figure 4 (Appendix C).
-Figure 2: Grid search results for the final return of USAC algorithm with Laplace distribution, averaged over three seeds, each with three evaluation episodes.
-The grid search explore various $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$ values, using SAC as the baseline.
-Figure 3: Grid search results for the estimation error (true - estimation) of USAC algorithm with Laplace distribution, averaged over three seeds, each with three evaluation episodes.
-The grid search explore various $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$ values, using SAC as the baseline.
-6 Discussion
-We introduce Utility Soft Actor-Critic (USAC), a novel off-policy actor-critic algorithm.
-The key idea is to balance the trade-off between pessimism and optimism through incorporating a utility function that captures the uncertainty in the critic due to limited access to the environment.
-This framework is flexible and can be adapted to different distributions, with our experiments primarily focusing on the Laplace distribution.
-Unlike OAC [7] and TOP [20], which also explore optimism/pessimism in actor-critic training, USAC provides a policy improvement guarantee.
-USAC allows independent control over the optimism and pessimism levels for the actor and critic.
-This decoupling enables a stable critic, necessary for accurate value function estimation, and an explorative actor, crucial for effective policy learning.
-We conducted a grid search experiment to explore these levels in training, revealing that being pessimistic in critic training while being either optimistic or pessimistic in actor training can be beneficial.
-However, it is also essential to consider the specific environment.
-Our work paves the way for further exploration in optimizing actor-critic algorithms with varying optimism and pessimism levels.
-Limitations and future work
-USAC’s primary limitation is its reliance on fixed optimism/pessimism parameters.
-An exciting direction for future research is to develop adaptive schemes where these parameters adjust dynamically throughout the learning process to better respond to the changing dynamics of the environment.
-There are several potential approaches to integrate automatic tuning.
-For example, these include: (i) adapting the automatic tuning scheme used for $\alpha$ in SAC [12, 13], (ii) employing bandit-based selection methods [20], or (iii) exploring continuous learning approaches using gradient descent [4].
-Another promising direction for future work is to incorporate higher moments of the critic distributions.
-By considering not only the mean and variance but also skewness, kurtosis, and heavy-tailed distributions, we can better capture the nuances of uncertainty in the critic’s value estimates.
-This approach could lead to more robust and flexible policy learning, accommodating a wider variety of distributional characteristics.
-Incorporating these higher-order moments may provide a richer framework for addressing the biases and variances in the value function estimation.
-Disclosure of funding
-This work was funded by the Novo Nordisk Foundation (NNF21OC0070621) and the Carlsberg Foundation (CF21-0250).
-Appendix A Proofs
-Proof of Proposition 1.
-Suppose $Q(s,a)\sim\textrm{Laplace}(\mu_{\mathcal{Q}}(s,a),b_{\mathcal{Q}}(s,a))$.
-Then, for any $(s,a)\in\mathcal{S}\times\mathcal{A}$, the expected utility function of $\mathcal{Q}$ on $(s,a)$ can be expressed as
+[…]
+Proof.
+Suppose $Q(s,a)\sim\textrm{Laplace}(\mu_{\mathcal{Q}}(s,a),b_{\mathcal{Q}}(s,a))$. Then, for any $(s,a)\in\mathcal{S}\times\mathcal{A}$, the expected utility function of $\mathcal{Q}$ on $(s,a)$ is
 $U^{\mathcal{Q}}_{\lambda}(s,a)=\mu_{\mathcal{Q}}(s,a)-\lambda^{-1}\log(1-\lambda^{2}b_{\mathcal{Q}}^{2}(s,a)),$
@@
 $U^{\mathcal{Q}}_{\kappa}(s,a)=\mu_{\mathcal{Q}}(s,a)-\sigma_{\mathcal{Q}}(s,a)\log(1-\kappa^{2})/\sqrt{2}\kappa.$
-At last, this expression can be further simplified using the definition of $g(\kappa)$ in 12.
-Appendix B USAC under gaussianity
+At last, this expression can be further simplified using the definition of $g(\kappa)$ in ˜12. ∎
+In Proposition˜2, we observe that by varying $\kappa$ in the small interval (1,1)(-1,1), $U^{\mathcal{Q}}_{\kappa}$ spans the entire spectrum of the $\mathcal{Q}$ distribution.
+Specifically, as $\lim_{\kappa\rightarrow-1}g(\kappa)\rightarrow$-$\infty$, $\lim_{\kappa\rightarrow 0}g(\kappa)\rightarrow 0$, and $\lim_{\kappa\rightarrow 1}g(\kappa)\rightarrow\infty$.
+Alternatively, one could also presume Gaussian-distributed $Q$-values.
 Proposition 3.
@@
 $U^{\mathcal{Q}}_{\lambda}(s,a)=\mu_{\mathcal{Q}}(s,a)+\lambda\sigma_{\mathcal{Q}}^{2}(s,a)/2\quad\text{for}\quad\lambda\in(-\infty,\infty).$ (13)
-Proof of Proposition 2.
-The proof follows by similar steps as in the proof of Proposition 1.
-Similarly, we can construct a USAC algorithm under Gaussian uncertainty, analogous to our approach under Laplace uncertainty as discussed in Section 4.3.
-By using the utility function defined in 16, we can incorporate specific $\lambda_{\text{critic}}$ and $\lambda_{\text{actor}}$ parameters for critic and actor training, respectively.
-This allows us to adjust the levels of optimism and pessimism tailored to the Gaussian-distributed $\mathcal{Q}$ values, thereby facilitating effective learning and exploration in our framework.
-Appendix C Experiments details
-Table 3: Properties of MuJoCo environments
-Environment Action dimension Observation dimension
-Ant-v4 8 27
-HalfCheetah-v4 6 17
-Hopper-v4 3 11
-Humanoid-v4 17 376
-Walker2d-v4 6 17
-Environments
-Our experiment is implemented in PyTorch [21, Version $2.1.0$].
-The experiments are conducted on 5 continuous control environments in the MuJoCo physics engine [30, 5].
-More detailed information about these environments are provided in Table 3.
-Table 4: Shared hyper-parameters
-Hyper-parameter Value
+Proof of Proposition˜3.
+The proof follows by similar steps as in the proof of Proposition˜2. ∎
+Now suppose the uncertainty follows a Laplace distribution.
+Following the algorithmic steps outlined in Section˜4.2, we can estimate the utility of the target critic distribution $\bar{\mathcal{Q}}$ by two target critics $\{\bar{Q}_{k}\}_{k=1,2}$, which are delayed versions of the critics $\{Q_{k}\}_{k=1,2}$.
+The utility function is given by $U^{\bar{\mathcal{Q}}}_{\kappa_{\text{critic}}}=\mu_{\bar{\mathcal{Q}}}+g(\kappa_{\text{critic}})\sigma_{\bar{\mathcal{Q}}}$ with utility parameter $\kappa_{\text{critic}}\in(-1,1)$, where
+$\mu_{\bar{\mathcal{Q}}}(s,a)=\frac{1}{2}(\bar{Q}_{1}(s,a)+\bar{Q}_{2}(s,a)),$ (14)
+and
+$\sigma_{\bar{\mathcal{Q}}}(s,a)=\frac{1}{2}\lvert\bar{Q}_{1}(s,a)-\bar{Q}_{2}(s,a)\rvert$. (15)
+Hence, the target value $y(s,a,r,s^{\prime})$ to train the critic becomes
+$r+\gamma[\mu_{\bar{\mathcal{Q}}}(s^{\prime},a^{\prime})+g(\kappa_{\text{critic}})\sigma_{\bar{\mathcal{Q}}}(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})],$ (16)
+with $a^{\prime}\sim\pi(\cdot|s^{\prime})$. Similarly, the actor learns its policy by
+$\operatorname*{arg\,max}_{\pi\in\Pi}\mathbb{E}_{a\sim\pi(\cdot|s)}[\mu_{\tilde{\mathcal{Q}}}(s,a)+g(\kappa_{\text{actor}})\sigma_{\tilde{\mathcal{Q}}}(s,a)-\alpha\log\pi(a|s)],$ (17)
+for $\kappa_{\text{actor}}\in(-1,1)$, where $\mu_{\tilde{\mathcal{Q}}}$ and $\sigma_{\tilde{\mathcal{Q}}}$ are estimated in similar manner as ˜14 and 15 on $\tilde{\mathcal{Q}}$ using the critics $\{Q_{k}\}_{k=1,2}$.
+The pseudo-code of this algorithm can be found in Algorithm˜1.
+We note that a smaller $\kappa$ value indicates more pessimism, while a larger $\kappa$ value indicates more optimism.
+By choosing $\kappa$ such that $g(\kappa)=-1$ (i.e., $\kappa\approx-0.83$), the utility in ˜11 simply reduces to the minimum of the two critics.
+Thus, choosing $\kappa_{\text{critic}}=\kappa_{\text{actor}}\approx-0.83$, we obtain the pessimism choices of SAC [14] and TD3 [12].
+On the other hand, by choosing $\kappa_{\text{critic}}\approx-0.83$ and a positive $\kappa_{\text{actor}}$, we have a pessimism critic and an optimistic actor as explored in OAC [9].
+Lastly, by letting a multi-armed bandit algorithm choose $\kappa_{\text{critic}}=\kappa_{\text{actor}}\in\{-0.916563,0\}$, we recover the pessimistic/optimistic design suggested by TOP [22].
+To recover the pessimistic case of the TOP algorithm, we can choose $\kappa=-0.916563$, which corresponds to having the function $g$ in ˜12 equal to $\sqrt{2}$.
+We note that by choosing 0.916563-0.916563, TOP considers a more pessimistic critic target than SAC, TD3, and OAC.
+5 Experiments
+We conduct experiments to answer two questions: (Q1) Can our utility-theoretic framework explain the previously observed relationships between the action-value biases and exploration optimism by two simple scalars, one for actor and one for the critic?
+(Q2) Can these scalars be steered to improve performance via a coarse grid-search?
+Throughout the experiments, we adopted the USAC version described in Algorithm˜1 that assumes Laplace-distributed critics as the representative implementation of our framework.
+We tested all models on the standard five continuous control benchmarks of the MuJoCo physics engine [33, 7] to facilitate the interpretation of our results and its comparability to previous studies.
+The implementation was done in PyTorch [23, Version $2.1.0$].
+We ran each experiment for one million time steps and tested each policy every $10,000$ steps over ten evaluation episodes.
+We repeated each experiment across five different seeds hard-set to numbers 1 to 5 for training.
+We chose seed numbers for evaluation as the training repetition number plus 100.
+We consider SAC [14], TD3 [12], OAC [9], and TOP [22] as baselines, as they represent the most powerful algorithms that take finding an optimal balance between optimism and pessimism as the main goal.
+[^1] For TOP, we use https://github.com/tedmoskovitz/TOP with the number of random actions set to 10,000.
+We modeled all the actor and critic networks as fully connected layers of $256$ hidden units that follow ReLU() activations.
+USAC, SAC, and OAC actors output parameters for a SquashedGaussian module via a linear layer of size $2\times d_{a}$, where the first outputs represent the mean and the second represent the variance.
+TD3 and TOP use deterministic policies with a linear layer of size $d_{a}$ followed by Tanh().
+Critic networks for USAC, SAC, OAC, and TD3 output scalars; TOP outputs $N=50$ quantiles.
+Here, $d_{s}$ and $d_{a}$ denote state and action dimensions.
+Note that although the paper of TOP describes a two-layer network, its code uses three layers; we modify it for consistency with other baselines.
+For SAC and USAC, we employ automatic entropy temperature tuning $\alpha$ [15].
+For OAC and TOP, we adopt the best parameters reported in their respective papers.
+The hyperparameters and network configurations shared by all models in our consideration are summarized in Table˜1.
+The baseline models do not have any additional hyperparameters required to be tuned.
+We provide a public implementation at https://github.com/adinlab/UtilitySoftActorCritic.
+Table 1: Shared hyperparameters
+Hyperparameter Value
 Evaluation episodes 10
@@
 $n$-step returns 1 step
-Replay ratio 1
+Replay ratio/update-to-data 1
 Replay buffer size 1,000,000
@@
 Mini-batch size $(n)$ 256
+Number of critics 2
 Actor/critic optimizer Adam [17]
@@
 Averaging parameter $(\tau)$ 5e-3
-Evaluation and seeds
-As mentioned in Table 4, we train our algorithms for one million time steps.
-Each experiment is repeated across 5 different seeds.
-We evaluate at every $10,000$ time step for ten evaluation episodes.
-For training, we use seeds 1 to 5, and for evaluation, we use seeds 101 to 105, which correspond to the training repetition number plus 100.
-Hyper-parameters
-The hyper-parameters and network configurations used in our experiments can be found in Table 4.
-For SAC and USAC, we learn the entropy temperature $\alpha$ during as outlined in Haarnoja et al.
-[13].
-For OAC [7] and TOP [20], we adopt the best parameters reported in their respective papers.
-Architectures and baselines
-The architectural specifications for all algorithms are presented in Table 5.
-Here, $d_{s}$ and $d_{a}$ represent the dimensions of the state and action spaces, respectively.
-The SquashedGaussian module features a Gaussian head, which uses the first $d_{a}$ inputs for mean calculation and the subsequent $d_{a}$ inputs for variance.
-This architecture is the standard one used in original implementations of SAC [12], TD3 [10], and OAC [7].
-However, while TOP [20] states that it employs a network architecture with $2$ hidden layers, its original implementation code actually considers $3$ hidden layers.
-[^6] We use the implementation provided by the authors for TOP, https://github.com/tedmoskovitz/TOP, where the "number of random actions" at the start of training updated to $10000$.
-This discrepancy prompted us to re-run TOP with a synchronized network such that it aligns with the other algorithms.
-In addition, the critic network in TOP follows the original implementation with a return shape that equals the number of quantiles (N-Quantiles), which is $50$.
-Table 5: Architecture details
-Actor network Critic network
-USAC
-SAC
-OAC
-TD3
-TOP
-USAC
-SAC
-OAC
-TD3
-TOP
-Linear($d_{s}$, 256)
-ReLU()
-Linear($d_{s}$ + $d_{a}$, 256)
-ReLU()
-Linear(256, 256)
-ReLU()
-Linear(256, 256)
-ReLU()
-Linear(256, 2×$\times d_{a}$, $\theta$)
-Linear(256, $d_{a}$) Linear(256, 1) Linear(256, N-Quantiles)
-SquashedGaussian(2×$\times d_{a}$, $d_{a}$) Tanh()
-Grid search of USAC
-In order to explore the dynamics of pessimism and optimism in USAC with Laplace distribution (Section 4.3), we conduct a grid search across five continuous control environments in MuJoCo (see e.g., environments specifications above).
-These grid search results are averaged over 3 seeds (seed 1 to 3), each with 3 evaluation episodes.
-This grid search investigated various values of $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$, using SAC as the baseline for comparison.
+Our framework can be implemented as a drop-in into any standard actor-critic pipeline as an alternative way to aggregate critic outputs into a single value.
+Hence, it does not cause a tangible computational overhead.
+Our representative USAC implementation adopts the twin-critic and single-actor architecture used in standard SAC, TD3, OAC, and TOP implementations.
+Its computation time is indistinguishable from a standard SAC implementation.
+Figure 2: The impact of the degree of actor optimism on the final return of USAC algorithm when Laplace distributed critics are assumed.
+The results are averaged over three seeds, each with three evaluation episodes.
+The experiments span various $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$ values, using SAC as the baseline.
+Figure 3: The impact of the degree of actor optimism on the estimation error (true-estimation) of USAC algorithm when Laplace distributed critics are assumed.
+The results are averaged over three seeds, each with three evaluation episodes.
+The experiments span various $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$ values, using SAC as the baseline.
+Figure 4: The impact of the degree of actor optimism on the learning speed of USAC algorithm when Laplace distributed critics are assumed, measured by the area under learning curve.
+The results are averaged over three seeds, each with three evaluation episodes.
+The experiments span various $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$ values, using SAC as the baseline.
+5.1 Question 1: Can USAC hyperparameters explain the optimism and pessimism dynamics?
+We investigate the performance profile of USAC with Laplace distributed critics (Section˜4.3) across different degrees of optimism and pessimism on five MuJoCo environments.
+We expect to replicate the findings of prior art [12, 9, 22], thereby verifying our claim that the trade-off problem can be reduced to tuning an interpretable hyperparameter.
+Results are averaged over three different seeds, comparing against SAC as a baseline, as it is a special case of USAC with $\kappa_{\text{critic}}=\kappa_{\text{actor}}=-0.831559$.
+This allows us to systematically analyze how different configurations of utility parameters affect learning behavior and overall returns.
 We examined critics ranging from pessimistic to semi-optimistic and actors from pessimistic to optimistic.
-Specifically, we used $\kappa_{\text{critic}}\in\{-0.831559,-0.5,-0.33\}$ [^7] A $\kappa=-0.831559$ corresponds to $g(\kappa)=1$, which means that the utility function in 11 simplifies into the minimum of the two critics.
-Hence, by setting $\kappa_{\text{critic}}\approx-0.83$, we align with the pessimistic choices of SAC [12], TD3 [10], and OAC [7]., and $\kappa_{\text{actor}}\in\{-0.99,-0.5,0,0.5,0.99\}$.
-The rationale for choosing more pessimistic (to semi-optimistic) $\kappa_{\text{critic}}$ values is to counteract the overestimation bias.
-Given that USAC stems from the framework of SAC, SAC was included in our figures for comparison.
-In Figures 2, 3, and 4, we present the results of our grid search, showcasing the final returns, estimation error, and area under the learning curves for various parameter combinations.
-The estimation error calculated by subtracting the estimated $Q$-value from the true (discounted) reward.
-Thus, positive values indicate underestimating, while negative value indicate overestimation.
-These findings provide insights into the dynamics of pessimism and optimism in our USAC algorithm.
-As mentioned in Section 5, they reveal that: (i) overestimation (negative estimation error) should always be avoided as it reduces performance, and (ii) underestimation (positive estimation error) does not necessarily harm performance.
-This suggests that a pessimistic approach should be adopted in critic training, while either a pessimistic or optimistic approach can be suitable for actor training, depending on the environment.
-The best combinations of $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$ in terms of final returns are summarized in Table 2.
-These optimal pairs highlight the importance of tailoring the levels of pessimism and optimism to the specific requirements of each environment to enhance overall performance.
-Figure 4: Grid search results for the area under the learning curve of USAC algorithm with Laplace distribution, averaged over three seeds, each with three evaluation episodes.
-The grid search explore various $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$ values, using SAC as the baseline.
-Learning curve and estimation error curves
-The learning curve and estimation error curves of our USAC algorithm and the baseline algorithms can be seen in Figures 5 and 6, which correspond to the experiments summarized in Table 1.
-In addition, the estimation error values for these experiments are provided inTable 6.
-For USAC, the legend corresponds to the default case where $\kappa_{\text{critic}}=\kappa_{\text{actor}}=-0.831559$.
-The USAC (best) legend uses the optimal $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ pairs (with respect to final return) as listed in Table 2.For the Humanoid-v4 environment, we found that the default parameter pair outperformed the best pair identified through grid search, e.g., see Figure 2.
-Therefore, we used the default pair for this environment in our results (Table 1).
-However, we conjecture that by expanding our (limited) grid search, a better pair may exist.
-Table 6: Estimation error of USAC and baselines.
+Specifically, we used $\kappa_{\text{critic}}\in\{-0.831559,-0.5,-0.33\}$ [^2] Recall that $\kappa=-0.831559$ corresponds to $g(\kappa)=1$, simplifying the utility function in 11 to the minimum of the two critics.
+Therefore, by setting $\kappa_{\text{critic}}\approx-0.83$, we align with the pessimistic choices of SAC [14], TD3 [12], and OAC [9]., and $\kappa_{\text{actor}}\in\{-0.99,-0.5,0,0.5,0.99\}$}.
+The rationale for choosing more pessimistic (to semi-optimistic) $\kappa_{\text{critic}}$ values is to counteract the overestimation bias commonly seen in off-policy reinforcement learning.
+We present results for three metrics that provide complementary information: (i) final return (Figure˜2) to assess the success of each configuration at solving the task at hand, (ii) estimation error (Figure˜3) to assess the explanatory value of our framework for further investigation, (iii) area under the learning curve (Figure˜4) to assess the learning speed of each configuration.
+The estimation error is calculated by subtracting the estimated $Q$-value from the true (discounted) reward.
+Thus, positive values indicate underestimating, while negative values indicate overestimation, both of which can impact learning efficiency and stability.
+Our results verify our hypothesis that the following key findings of the prior art can indeed be explained by a single hyperparameter value:
+  • Overestimation (i.e., negative estimation error) should always be avoided as it reduces performance, as visible from the strong correlation between high estimation error in Figure˜3 and low final return in Figure˜2.
+  • Underestimation (i.e., positive estimation error) does not necessarily harm performance.
+This outcome reaffirms the central claims reported in the TD3 paper [12], which yielded the wide adoption of the min-clipping method.
+We also conclude from the results that a pessimistic approach should be adopted in critic training, whereas in actor training, either a pessimistic or optimistic approach can be suitable depending on the environment and the exploration-exploitation balance required.
+This outcome reaffirms the prior findings [12, 9, 22], supporting our hypothesis about the possibility of expressing the optimism and pessimism dynamics on a simple and interpretable axis.
+Table 2: The best pair of hyperparameters $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ for USAC with respect to final returns; see for example Figure˜2.
+Environment $\kappa_{\text{critic}}$ $\kappa_{\text{actor}}$
+Ant-v4 0.831559-0.831559 0.99-0.99
+HalfCheetah-v4 0.33-0.33 0.50-0.50
+Hopper-v4 0.831559-0.831559 $0.50$
+Humanoid-v4 0.831559-0.831559 0.831559-0.831559
+Walker2d-v4 0.831559-0.831559 $0.0$
+5.2 Question 2: Can the hyperpararameters be effortlessly tuned to improve performance?
+From Figures˜2, 3, and 4, we can extract the best combinations of $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ per environment with respect to final return and treat them as the outcome of a search on a coarse grid of hyperparameters.
+These best values are listed in Table˜2.
+The difference between the best values across the environments highlights the environment-specific nature of optimism and pessimism and also demonstrates the success of our framework at detecting these differences.
+Table 3: Final return on MuJoCo environments trained with 1M time steps, averaged over 5 seeds.
+The best algorithms are highlighted in bold.
+$\pm$ corresponds to the standard deviation across repetitions.
+$(\kappa_{\text{critic}},\kappa_{\text{actor}})$ for USAC are listed in Table˜2.
 Environment USAC (ours) SAC TD3 OAC TOP
-Default $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ Best $(\kappa_{\text{critic}},\kappa_{\text{actor}})$
-Ant-v4 $80.66$ $91.84$ $75.33$ $196.77$ $205.49$ $98.99$
-HalfCheetah-v4 $221.21$ $146.07$ $222.93$ $216.54$ $190.28$ $81.97$
-Hopper-v4 $12.46$ $18.51$ $12.73$ $24.82$ $25.05$ $11.73$
-Humanoid-v4 $40.24$ $40.24^{\star}$ $42.78$ $31.89$ $52.66$ $24.02$
-Walker2d-v4 $42.50$ $44.87$ $36.65$ $50.88$ $52.55$ $34.34$
+Ant-v4 $\boldsymbol{5158\pm 424}$ $4756\pm 773$ $4091\pm 130$ $4177\pm 802$ $4334\pm 1276$
+HalfCheetah-v4 $\boldsymbol{11736\pm 297}$ $10763\pm 891$ $10570\pm 801$ $8684\pm 1671$ $7311\pm 3074$
+Hopper-v4 $\boldsymbol{3442\pm 126}$ $3185\pm 380$ $1986\pm 1114$ $293\pm 112 33$$369\pm 154$
+Humanoid-v4 $\boldsymbol{5602\pm 210}$ $5503\pm 198$ $5149\pm 221$ $5390\pm 121$ $5332\pm 445$
+Walker2d-v4 $\boldsymbol{4530\pm 708}$ $3757\pm 1069$ $4369\pm 594$ $3467\pm 1193$ $4317\pm 631$
 Figure 5: Learning curves of USAC and baselines. Solid curves depict the average return across evaluation episodes, while the shaded areas represent the standard deviation.
-Figure 6: Estimation error curves of USAC and baselines.
-Solid lines indicate the estimated value function by the critic, while dashed lines show the averaged episodic discounted true return.
-The shaded areas around the solid and dashed lines represent the standard deviation from the average estimated value and the average true reward, respectively.
-Computational cost
-Our USAC algorithm uses twin critics and a single actor network as for SAC, TD3, OAC, and TOP.
-This design choice eliminates the need for any additional computational steps and is no reliance on repetitive sampling, pre-training, or iterative approximation procedures.
-Consequently, it can be conclude that its computational characteristics closely match those of other widely used actor-critic algorithms and ours comparing baselines reported.
-The code for our algorithm implementation is also available alongside the paper.
+We give a performance comparison of the grid-optimized USAC against the baseline models in Table˜3.
+The superior performance of USAC supports our hypothesis that the hyperparameters introduced by our framework also have an instrumental value, as they can be used to improve actor-critic algorithms.
+It is also remarkable that USAC incurs a smaller standard deviation than the baseline models across repetitions in the far majority of the cases, indicating an improved model stability during learning.
+The learning curve of the grid-optimized USAC algorithm is given in Figure˜5 together with the baseline algorithm.
+The results suggest a consistent performance improvement that spans the whole learning regime.
+In applications where online learning with minimum regret is more important than maximizing the final reward, such as interactive learning applications, a similar grid-search may also be performed with respect to the area under curve.
+While we omit a detailed reporting of such an analysis for brevity, one can see from Figure˜4 that USAC will be similarly competitive in that scenario.
+6 Discussion and limitations
+We introduced a novel algorithmic framework to trade pessimistic policy evaluation and pessimistic/optimistic policy search in off-policy deep actor-critic pipelines.
+The key idea is to balance optimism and pessimism through incorporating a utility function that captures the uncertainty in the critic due to limited access to the environment into a single and interpretable hyperparameter.
+This USAC framework is flexible and can be adapted to different distributions, with our experiments primarily focusing on the Laplace distribution.
+Our USAC framework allows independent control over the optimism and pessimism levels for the actor and critic.
+This decoupling enables a stable critic, necessary for accurate value function estimation, and an explorative actor, crucial for effective policy learning.
+Furthermore, USAC allows tuning the degree of optimism/pessimism via a single hyperparameter that interpretably encodes full pessimism as -1 and full optimism as +1 and all intermediate possibilities as values in between.
+Our experiments revealed that being pessimistic in critic training while being either optimistic or pessimistic in actor training can be beneficial.
+However, it is also essential to consider the specific environment.
+We also found that a search on a coarse grid of our new hyperparameters brings consistent performance gains across environments with diverse characteristics.
+The primary limitation of USAC is that it does not prescribe any insights for on-the-fly tuning of its hyperparameters.
+An exciting direction for future research is to develop adaptive schemes where these parameters adjust dynamically throughout the learning process to better respond to the changing dynamics of the environment.
+There are several potential approaches to integrate automatic tuning.
+For example, (i) adapting the automatic tuning scheme used for $\alpha$ in SAC [14, 15], (ii) employing bandit-based selection methods [22], or (iii) exploring continuous learning approaches using gradient descent [6].
+Another limitation is our evaluation scope.
+We focus on the five most used MuJoCo environments [33], where the studied phenomenon and the impact of our solution are clearly visible, ensuring comparability with the key baselines we consider.
+Broader validation across additional domains would further demonstrate generality, which we plan to explore in future work.
+Our findings could also be improved by incorporating higher moments of critic distributions (mean, variance, skewness, kurtosis) to better capture uncertainty in value estimates.
+Lastly, extending from critic pairs to ensembles may offer faster training at the cost of increased computational power.
+Acknowledgements
+This work was funded by the Novo Nordisk Foundation (NNF21OC0070621) and the Carlsberg Foundation (CF21-0250).